Some ideas for improving L^{A}TEX

David Rhead University of Nottingham

JANET: d.rhead@uk.ac.nottingham.ccc.vme

July 1991

Contents

Preface iv

1 General 1 1.1 Assumptions about structure . . . . . . . . . . . . . . . . . . 1 1.2 Use of counts . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3 Ragged right . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.4 File names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2 Preliminary pages 8 2.1 Title-page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2 Those pages without printed numbers . . . . . . . . . . . . . 8

3 Floats 9 3.1 Confusion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.2 Captions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

3.2.1 L^{A}TEX 2.09’s practice . . . . . . . . . . . . . . . . . . . 10 3.2.2 Publishing practice . . . . . . . . . . . . . . . . . . . . 11 3.2.3 L^{A}TEX 3.0 . . . . . . . . . . . . . . . . . . . . . . . . . 12

3.3 Landscape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.4 E-mail following “L^{A}TEX3 and you” meeting . . . . . . . . . . 13

4 Tables 22 4.1 Decimals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.2 Rules, i.e. lines . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.3 Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.4 An e-mail addition . . . . . . . . . . . . . . . . . . . . . . . . 24 5 Lists of references; bibliographies 26

5.1 Title . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 5.2 Schemes supported . . . . . . . . . . . . . . . . . . . . . . . . 27 5.3 Consequences for BibTEX . . . . . . . . . . . . . . . . . . . . 29 5.4 Subdivisions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 5.5 Note . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

6 “Over-long” headings 30

7 Page headers 32

8 Sections 33

9 Lists 39 9.1 Within paragraphs . . . . . . . . . . . . . . . . . . . . . . . . 39 9.2 An interrupted enumerate? . . . . . . . . . . . . . . . . . . . 40

10 verbatim  41

10.1 Vertical space . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 10.2 Horizontal space . . . . . . . . . . . . . . . . . . . . . . . . . 42

11 Borrowing some margin 43

A Some e-mail comments about structure 45

B Support for accurate page-selection from dvi files 63 B.1 Objectives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 B.2 A scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 B.3 Some e-mail comments . . . . . . . . . . . . . . . . . . . . . . 67

C Some e-mail comments about treatment of wide objects 69

D Citations and reference-lists 78 D.1 Some e-mail comments . . . . . . . . . . . . . . . . . . . . . . 78 D.2 My e-mailed “paper” . . . . . . . . . . . . . . . . . . . . . . . 88

D.2.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . 88 D.2.2 Conventions to be supported . . . . . . . . . . . . . . 89 D.2.3 Deficiencies in L^{A}TEX 2.09 . . . . . . . . . . . . . . . . 90 D.2.4 Introduction of new facilities . . . . . . . . . . . . . . 93 D.2.5 Division of labour . . . . . . . . . . . . . . . . . . . . 93 D.2.6 L^{A}TEX 3.0: A possible user interface? . . . . . . . . . . 95 D.2.7 Details of reference-lists . . . . . . . . . . . . . . . . . 99 D.2.8 Conclusion . . . . . . . . . . . . . . . . . . . . . . . . 102 Appendix: Interface with BibTEX . . . . . . . . . . . . . . . . 103

D.3 Further e-mail comments . . . . . . . . . . . . . . . . . . . . 105

E Some e-mail comments on standard styles 109

F Could L^{A}TEX do more for chemists? 121 F.1 Motivation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 F.2 Typesetting and artwork . . . . . . . . . . . . . . . . . . . . . 122 F.3 Desirable facilities for chemists in the new L^{A}TEX . . . . . . . 123

F.3.1 Environments . . . . . . . . . . . . . . . . . . . . . . . 123 F.3.2 Commands . . . . . . . . . . . . . . . . . . . . . . . . 124

F.4 Work involved . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 F.5 Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . 126 F.6 Some e-mail comments . . . . . . . . . . . . . . . . . . . . . . 126

Preface

Frank Mittlebach suggested that I might produce some material for consideration by L^{A}TEX-people at Dedham. This document is the result.

My involvement with L^{A}TEX 3.0 [1, 2] started early in 1990 when I heard Chris Rowley announce that opinions were invited about the facilities that should be provided in the new L^{A}TEX. Accordingly, I submitted a “wish list”, and have since submitted various additional ideas.

The present document is an attempt to put these wishes etc. together in one place. It has been put together in a hurry from various sources. Generally the main text is based on older material than the appendices: you may find inconsistencies of detail that I haven’t had time to resolve. (Frank only gave me two weeks notice!) I’ve included some extracts from e-mail in which people were making comments about similar topics. (My “extracting algorithm” has been an uneasy compromise between thinking that I shouldn’t confine the document to my own views, and wondering whether people would want remarks they made in one context reproduced in another context. I apologise to anyone whose views I’ve omitted/included when I should have included/omitted them.) I realize that the result is overlong and repetitive, but I don’t have time to think how to cut it down and re-organise it.

In my case, most “wishes” are just particular cases of general wish “that L^{A}TEX would make it easy to produce documents that conform to the conventions that are usual in traditional publishing”, since most of the end-users who tell me “I don’t like the way L^{A}TEX does this” are, I think, really saying “I’d just like L^{A}TEX to do things like normal publishers do them” (either because they want to pass the result on to a publisher, or simply because traditional conventions “look right” to them).

I’ve generally tried to determine “traditional publishing practice” by consulting the manuals to which a publisher’s copy-editors, designers, and typesetting staff might refer. I assume that the Chicago Manual of Style [3], Butcher’s Copy-editing, [4], McLean’s Manual of Typography [5], Williamson’s Methods of Book Design [6] and Hart’s Rules [7] are representative of such manuals. Since I assume that L^{A}TEX is, and always will be, intended for producing formally structured documents rather than for implementing “gridbased designs”, I presume that it is unrealistic to try to support more than a selection of the features described in books (such as [8] and [9]) that are aimed at the “design for desktop publishing” market. However, the classes of document that are of interest to the Text Encoding Initiative [10] seem similar to the classes for which L^{A}TEX is intended, so I assume that the Initiative’s views on structure are relevant (although their objectives are broader than just typesetting).

I’m very aware that it’s easier “to wish” than “to grant wishes”. viii

Chapter 1

General

1.1 Assumptions about structure

L^{A}TEX 2.09’s assumptions about the structure of a book or report seem at odds with the usual assumptions in publishing practice. Consequently, the author often ends up doing things that are really the designer’s responsibility.

In publishing practice, a book-like document would have the following major divisions:

• “preliminary pages” or “front matter”, containing table of contents, acknowledgements, etc. These pages are usually numbered in roman.

• main text, containing the chapters.

• “end pages” or “back matter”, containing appendices, bibliography, index, etc.

See [3, p. 4–5], [4, p. 119, 165–166], [6, ch. 8], [5, ch. 10], [9, p. 157–161] and [10, p. 73–77].

Thus, for “front matter” and “back matter”, the divergence between the real structure and that assumed by L^{A}TEX 2.09 forces the author to think visually rather than logically. The author gets involved with:

• \chapter* plus \addcontentsline for most “top level” units in the front and back matter.

   • having to put \pagenumbering{roman} and \pagenumbering{arabic}

around the “front matter”.

\documentstyle[11pt]{report}
\begin{document}       %  A single-sided project report.
\pagenumbering{roman}  %  Table of contents and other preliminary pages are
\setcounter{page}{2}   %  to be numbered in roman.  If numbering notionally
                       %  starts with the title-page (which is being typeset
                       %  separately), actual numbering should start at "ii"
                       %  on the table of contents.
\tableofcontents
\include{other-front-matter}
\clearpage             %  Switch to arabic page-numbering for the rest of
\pagenumbering{arabic} %  the report.
\include{a-chapter}
\include{another-chapter}
\appendix              %  Arrange that \chapter will actually give units
                       %  entitled "Appendix ...".
\include{an-appendix}
\include{another-appendix}
\include{other-back-matter}
\end{document}

Figure 1.1: A L^{A}TEX 2.09 root file

\chapter*{Acknowledgements}
\addcontentsline{toc}{chapter}{Acknowledgements}
Text text text text text text text text text text text text text text text text.
Text text text text text text text text text text text text text text text text.
Text text text text text text text text text text text text text text text text.
Text text text text text text text text text text text text text text text text.
\chapter*{Preface}
\addcontentsline{toc}{chapter}{Preface}
Text text text text text text text text text text text text text text text text.
Text text text text text text text text text text text text text text text text.
Text text text text text text text text text text text text text text text text.
Text text text text text text text text text text text text text text text text.

               Figure 1.2: other-front-matter.tex

\chapter*{Glossary}
\addcontentsline{toc}{chapter}{Glossary}
Text text text text text text text text text text text text text text text text.
Text text text text text text text text text text text text text text text text.
Text text text text text text text text text text text text text text text text.
Text text text text text text text text text text text text text text text text.
\begin{thebibliography}{width of label}
\addcontentsline{toc}{chapter}{Bibliography}
\bibitem[Item label]{a-book} Author author author.
      {\em Title title title title title.\/}
      Publisher, etc.
\bibitem[Another label]{another-book}
      Another author another author.
      {\em Another title another title another title.\/}
      Another publisher, etc.
\end{thebibliography}

                Figure 1.3: other-back-matter.tex

See Figures 1.1, 1.2 and 1.3.

I think it would be better if L^{A}TEX 3.0 worked more in terms of the concepts used in publishing practice. The author might, for example, create an input file of the form shown in Figure 1.4. Then details such as romannumbering, table-of-contents entry, etc. could be dealt with in the style file (as questions of “design”) without the author being involved. Although the author would have to learn a bit of publishing jargon (“front matter”, etc.), this might at least help him/her to understand the language used in books about typesetting and by publishers. Most authors would no longer need to

learn about the the \chapter*, \addcontentsline and \pagenumbering

commands. (The \chapter* and \addcontentsline commands might appear in style files, where there could be a distinction between the table of contents, which does not need an \addcontentsline, and the other front and back matter units that, for many designs, do need an \addcontentsline.)

Obviously decisions will need to be taken about the terminology, commands and environments to be used for markup schemes. Figure 1.4 gives the basis for one scheme. The Text Encoding Initiative’s terminology [10, p. 73–77] is another option. Presumably another scheme could be produced by studying the Association of American Publishers’ recommendations about SGML markup (although I haven’t yet been able to study them myself) [11, 12, 13, 14, 15]. There would presumbly be advantages in some compatibility with conventions used outside the L^{A}TEX community.

Appendix A contains some extracts from e-mail discussions about such matters.

1.2 Use of counts

Although TEX provides 10 \count registers [16, p. 119] for page-identification, neither the “plain” macros nor the L^{A}TEX 2.09 standard styles use anything but \count0. From the point-of-view of a L^{A}TEX user, this can create situations in which page-selection is difficult at the previewing/printing stage.

For example:

• in the L^{A}TEX 2.09 report and book styles, both (roman-numbered) page iv in the “front matter”, and (arabic-numbered) page 4 in chapter 1 have a \count0 value of 4 in the dvi file, so it is difficult to instruct a dvi-file processing command to print page 4 rather than page iv.

• if a style-file for a document that numbers pages “within chapters”

\documentstyle ...
\begin{document}                                 % With this information, the
% style file should be able
\begin{frontmatter}                              % to work out where to
% start/stop roman numbers,
\tableofcontents                           % and that "Preface" is to
% appear in the table-of-
\toplevel{Preface}  %  \chapter* is a      % contents but "Table of
%  bad name for a      % contents" is not to appear
...                 %  top-level unit      % in itself.  I.e. author
%  that's NOT a        % specifies structure; design
\end{frontmatter}         %  chapter"            % decisions go in style file.
\begin{maintext}                                 % With this information, a
% style file could, if desired,
\chapter{A chapter}                        % put the word "Chapter" in
% in the toc file at the start
...                                        % of the list of chapters in
% the table of contents.
\chapter{Another chapter}                  % Similarly, with the
% information below, the style
...                                        % file could put "Appendix"
% in the toc file above the
\end{maintext}                                   % list of appendices.
\begin{backmatter}                               % It would be nice if,
% with this information
\begin{appendices}                         % and the aux file, the
% style file could work
\toplevel{An appendix}               % out whether there is
% only one appendix
...                                  % and, if so, give a
% heading of "Appendix"
\toplevel{Another appendix}          % rather than "Appendix A".
% If not, perhaps the
...                                  % user should have the
% choice of \begin{appendix}
\end{appendices}                           % of \begin{appendices} so
% as to avoid "Appendix A"
\toplevel{Glossary}                        % when there is no
% "Appendix B".
...
% Within frontmatter and
\toplevel{Index}                           % backmatter, a 3.0 style
% file might treat
...                                        % \toplevel like 2.09's
% {\chapter*
\toplevel{References}                      % + \addcontentsline}.
...
\end{backmatter}
\end{document}

Figure 1.4: A better way to specify the units in a book? follows the precedent of the standard styles and just uses \count0, page-selection will be difficult because there will often be several pages with the same \count0 value. If a dvi file represents 10 chapters, it will be difficult to select the first page of chapter 9, because dvi-processing software is not usually designed for tasks such as “find the 9th time that \count0 has the value 1”.

I think that the L^{A}TEX 3.0 standard styles should use the \counts in such a way that pages can be selected unambiguously at the dvi-file processing stage. Appendix B suggests some schemes that might achieve this objective.

1.3 Ragged right

People may wish to implement “ragged right” designs. They may want TEX’s mathematical typesetting capability and the help that L^{A}TEX can give with section-numbering, cross-referencing, etc., but not want any of the rightjustification that TEX supplies by default. This seems respectable, since there are books on design that advocate “ragged right” [9, p. 19].

Unfortunately, to implement a “ragged right” design in L^{A}TEX 2.09, it seems necessary to make individual changes to any definitions in a style file that involves “paragraph makeup”. For example, it is necessary to specify “ragged right” for main text, footnotes, captions and list-elements separately.

However, if the “main text” is to be typeset “ragged right”, it is most likely that “ragged right” will be required everywhere else, too. Could there be some easy mechanism for specifying, in a style file, that “everything is to be ragged right”? Better still, could there be some easy way of choosing, in a style file, to have everything as what White [9, p. 21] calls “rough rag”, “tight rag” or “minimal rag”?

1.4 File names

Might it be worth introducing a naming convention such as:

• extension sty for a style file

• extension opt for a style-option file?

It may help people to find their way round archives such as that at Aston.

An e-mail comment

From:     Don Hosek <DHOSEK@EDU.CLAREMONT.HMCVAX>
Date:     Fri, 1 Feb 91 21:54:00 PST
...
I'd like to see LaTeX 3.0 use different extensions for style
files and options. My recommendations would be style files have
extensions of .dst and options .dop (.opt is too commonly used
for other purposes)  ...  part of the
advantage gained by the new extensions would be the emphasis that
style files will probably need mods to work under the new
LaTeX.
...

Preliminary pages

2.1 Title-page

In L^{A}TEX 2.09, titlepage does a \setcounter{page}{0}. Unfortunately, when twoside is used, this setting leads to the titlepage having the margins appropriate for a left-hand page, when it should really have the margins for a right-hand page.

Since \thispagestyle{empty} is used, the actual number makes no difference in titlepage. Presumably the intention is to ensure that the following page is, by default, numbered “1”.

Could some way be found of ensuring that a title-page is given the margins for a right-hand page even when twoside is in effect?

This might well come about anyway if the markup scheme for reports and books is reconsidered (as suggested in section 1.1). If not, perhaps “0” should be treated as “odd” for the purpose of margins.

2.2 Those pages without printed numbers

It seems as though, in publishing practice ([3, p. 4], [4, p. 120]) the first 4 or 6 “preliminary pages” of a book will probably be un-numbered, the title-page being just one of these. If the user is expected to use such an environment for pages other than the title-page [1, p. 162], it might be worth thinking of a more appropriate name (unnumbered, perhaps). On the other hand, if the markup scheme for reports and books is reconsidered (as suggested in section 1.1), the situations in which people are expected to use titlepage for pages other than the title-page might be eliminated anyway.

Floats

3.1 Confusion

The figure and table environments arrange for:

• captions to be numbered

• material to be floated.

Our L^{A}TEX users seem to want the caption-numbering, but to get confused by the floating. Typically, they go \begin{table}[h] without reading the “small print” [1, p. 176–178], and then come to our Advisory Service asking why the table hasn’t appeared “here” and why they’re being told

“Too many unprocessed floats”.

Admittedly, much of the problem is that (a) they haven’t studied printed books to see what “the professionals” do with figures and tables (b) there is a typewriter tradition of expecting things to stay where they are put (c) they get a half-truth “from a friend”, and don’t study the manual to get the whole truth.

But, I think part of the problem is attributable to:

• the environment names. “table” sounds as if it’s “what you must use for a table”

• the dual role. People may choose table and figure in order to get numbered captions: there isn’t another off-the-shelf way to get them. They may see the floating as an unwanted side-effect and think (wrongly) that [h] will turn the effect off.

In L^{A}TEX 3.0, could environments for “floating” be given names that make it obvious that their main purpose is to float things?

For example, suppose that someone had to go \begin{floattable} or \begin{float}{table}. Even if they found out about this “from a friend”, they would stand a good chance of realising that they are asking L^{A}TEX to float something (and get a hint that h is “allowing a null float when aesthetically possible” rather than “forcing the table here under all circumstances”).

I’m not sure whether to be optimistic or pessimistic about what would happen if floating environments were given names that made it absolutely clear that their role is to float things. Being optimistic, I might think that users would then be clear what to expect, would be pleased that L^{A}TEX would make their document look like a professionally produced book, and would not use the environments in such a way as to give “Too many unprocessed floats” over-frequently. Being pessimistic, I wonder if it would be worth providing one or other of:

• environments that just put a table or figure “here” (with spacing, caption, etc., as for the “floats”), and give an “Underfull vbox” if the table/figure can’t be fitted on the current page and has to go on the next one (leaving the current page underfull). For example,

     \begin{fixedtable} or \begin{fixed}{table}.

• a command (or commands) to produce captions in “ordinary text” that look the same as (and use the same counters as) the captions produced in the floating environments. For example, \caption{table}{...} or

     \tablecaption{...}.

3.2 Captions

3.2.1 L^{A}TEX 2.09’s practice

The L^{A}TEX 2.09 manual [1] seems vague about whether the spacing around \caption is designed to go under or over a table/figure. Looking at the standard doc files suggests that the same code is used for both figures and tables, and that the resulting caption is intended to go under the figure or table.

However, the 2.09 manual points out that “there can be several captions in a figure or table environment”.

3.2.2 Publishing practice

In publishing practice, it seems respectable to have a combination of:

• table captions above the table

• figure captions below the figure.

Perhaps this is because tables may have notes underneath, and may be split over several pages. See [3, p. 266,308], [4, p. 136,160] and [7, p. 158,161].

Within such floats, it seems normal [3, ch. 11,12], [4, p. 164], [7, figs. 2,5] to have the following elements:

• for tables: the title (and possibly a subtitle), the table itself, source notes, general notes, specific notes, probability levels^{1}

• for figures: the figure itself, the caption, the legend.

I haven’t been able to find any sensible examples of floats with more than one caption. Hart has some examples of figures side-by-side [7, figs. 6,7,8], but this is really a case of figures ending up together after floating, rather than having to be kept together while floated. Butcher mentions situations where “two tables should appear in the same opening, so that they can be compared” [4, p. 161], but some mechanism other than that provided by L^{A}TEX 2.09 would be necessary to achieve this.

Overall, the vagueness about how the author and the designer are to use \caption seems likely to:

• make it difficult for the designer to implement decisions about the positioning, spacing and fonts for

– table titles and notes

– figure captions and legends

• leave the author doing visual design: experimenting until things “look right”.

1A statistician colleague tells me that statisticians now prefer “confidence intervals”.

3.2.3 L^{A}TEX 3.0

It seems to me that the advantage of having the usual L^{A}TEX “separation of the designer’s and author’s roles” within the body of a float would far outweigh any flexibility that “several captions within a float” allows. (If it is desired to allow two narrow floats side-by-side, this should be done by letting the user tell L^{A}TEX about the width of the float. If it is desired to cater for rare situations where “two tables should appear in the same opening”, this should be done by arranging some “magnetic attraction” between the two—or by letting the user fiddle with the tex file.)

I suggest that:

• notice be given that “more than one caption within a float” will not be supported after L^{A}TEX 2.09

• at L^{A}TEX 3.0, ideally

– the successor to table should require the user to specify the block to be floated in terms of the logical elements “title” (plus possible “subtitle”), “table itself”, “source notes”, “general notes”, “specific notes”, “probability levels”

– the successor to figure should require the user to specify the block to be floated in terms of the logical elements “figure itself”, “caption”, “legend”

– the positioning, typeface, etc. to be used for the logical elements should be under the control of the document-designer via the style file. A designer should be able to implement designs [3, 4, 7] that have: table-title at the top of a floated “table block”, with tablenotes at the bottom of the block; “caption” and “legend” (in that order) at the bottom of a floated “figure block”. Equally, a designer should be able to implement other conventions for layout of the “logical elements” within floated blocks.

3.3 Landscape

The precise details of what one does to place an object “landscape” will depend on the \specials available for a particular printer.

It may be unwise to define official environments, since they would not be available to everyone that has L^{A}TEX 3.0. (But if you want to define a standard “optional extension”, you might adopt definitions like those shown in Figure 3.1, which are based on code sent to UKTeX by Sebastian Rahtz.)

Nevertheless, people are going to want to define environments that place objects “landscape”. (Perhaps designers will want to provide such environments as part of an overall design.) It might be possible to help such people by providing the ability to define:

• a float that must appear on a “page of floats” on which it is the only float

• a \thispagestyle that takes effect after floating. Although there doesn’t seem to be a consensus between the gurus about whether one should print headings on pages with full-page tables ([3, p. 343], [7, p. 157]), those designers who do want to suppress the headings will need a suitable mechanism.

See also Appendix C.

3.4 E-mail following “L^{A}TEX3 and you” meeting

From:     David Rhead ...
Date:     18 Feb 91 14:52:55
BACKGROUND
There was a meeting called "LaTeX3 and you" on 14th Feb.  It had
been convened as part of Frank and Rainer's visit to London.
One of the topics discussed at the meeting was the idea of catering for
"floats" whose width may be more (e.g. if the text is being typeset in
2 columns) or less (e.g. if 2 narrow illustrations will fit horizontally next
to each other) than the width of the text-column.  Some notes of ideas
were distributed, but I'm afraid that I don't have time to type them in
here.
HART'S RULES
Hart and his successors at Oxford University Press have attempted to
write down some of their experience of how to deal with such things:
      "Hart's Rules [for Compositors and Readers at the University
      Press Oxford]", 39th edition, Oxford University Press, 1983.
      ISBN 0-19-212983-X.
I'll summarise Hart's recommendations, but first I'll explain the

%  Declare a storage bin for use by environments that deal with wide things.
\newsavebox{\widethingstore}
%  Deal with very wide objects as follows:
%  - typeset the object into a storage bin
%  - reserve a page (by giving LaTeX a portrait object that will occupy a
%    whole page)
%  - bring the object out of the storage bin into a region of the page that
%    dvitops will rotate.  Position the object so that, when rotated, its
%    vertical centre is aligned with the horizontal centre of the usual
%    portrait text-column.
\newlength{\height}  \newlength{\excess}
\newenvironment{landscapetable}{%
\begin{table}[p]
\savebox{\widethingstore}\bgroup\begin{minipage}{237mm}}{%
%
\end{minipage}\egroup
%  Will the object be too high for the landscape page?
\height=\dp\widethingstore  \advance\height by \ht\widethingstore
\excess=\height
\ifdim\height>127mm
    \advance\excess by -127mm
    \immediate\write16{Warning: Your landscapetable is \the\excess}
    \immediate\write16{higher than the normal landscape text-area.}
\fi
\vspace*{241mm}  %  A portrait object that will occupy a whole page.
\hspace*{64mm}   %  Align the vertical centre of the landscape object
                 %  with the horizontal centre of usual portrait text.
\special{dvitops: origin landscapetext}
\special{dvitops: begin  landscapetext}
\dp\widethingstore=0pt    %  Pretend that the object has no size,
\wd\widethingstore=0pt    %  otherwise it ends up in the wrong place.
\ht\widethingstore=0pt
\usebox{\widethingstore}  %  Bring the object out of its bin to
\special{dvitops: end}    %  be printed in a landscape orientation.
\special{dvitops: rotate landscapetext -90}
\end{table}
}
\newenvironment{landscapefigure}{%
\begin{figure}[p]
...
}

Figure 3.1: Prototype definitions of landscapetable and landscapefigure

terminology (in case anyone else reads Hart's Rules and is as
puzzled as I was at first):
Opening - A 2-page spread.
Block - "a relief printing surface from which line-art, half-tones,
  etc. can be printed".  The term comes from the wood blocks in
  which illustrations used to be cut.  (I guess that the analogue
  these days would be the "bounding box" for a file of encapsulated
  PostScript.)
Underline - A line [of text] underneath a block (NOT an \hrule underneath
  a line of text).  Thus, I think Hart's Rules uses "underline" to
  mean what the Chicago Manual of Style would call "caption plus
  legend" and what LaTeX 2.09 might call a \caption.
  The "underline" might contain an acknowledgement of a source.
Illustration - A block plus its underline.  I.e., in LaTeX 2.09 terms,
  the stuff that might go in \begin{figure} ... \end{figure}.
Here's my interpretation of Hart's Rules about how one should try to
float illustrations [where (a), (b), etc. are Hart's letters]:
(a) Keep the illustration on the same 2-page spread as the reference to
it as far as possible.
(b) If an illustration takes half-a-page (vertically) or less, it
should be placed slightly above the centre of the page, with
text above and below.  (I.e., in LaTeX 2.09 terms, there should
be an alternative to [t], [b] and [h] that gives "slightly
above the centre of the page".  It might need to try to avoid
splitting paragraphs:  see notes about Williamson, below.)
(c) If an illustration takes more than half a page, put it at the top
or bottom of a page (preferably the top), but don't ever put less
than 5 lines of text on a page.
(d) For a pair of full-width illustrations that have to appear in the same
2-page spread:  if the depth allows, they can  be put together on a page;
if not, they can be placed as for (b) or (c).  Hart's Rules gives
an example (figure 2) that has the following arrangement for a 2-page
spread:
        ILLUSTRATION      TEXT
           TEXT       ILLUSTRATION
(e) If 3 full-width illustrations have to be fitted onto a 2-page spread,
put 2 on one page and the third on the other with some text (minimum
4 lines) on the other.
(f) Hart's (f) deals with "running text around a narrow illustration",
which is probably outside the scope of LaTeX 3.0.  However, it
does mention typesetting underlines (i.e. caption + legend, etc.)
"to the width of the illustration", which may be a requirement
to bear in mind.
(g) If a "block" (i.e., piece of artwork) is narrow, move the
underline (i.e., caption + legend) to the side of the block,
so as to avoid having to run text around an illustration (i.e.,

    narrow artwork with narrow caption underneath).
    If you do this, the alignment of the caption is to the top/bottom
    of the block depending on whether the illustration is at the
    top/bottom of the page.
(h) If the width of an an illustration is less than full-width but greater
    than full-width/2, centre it horizontally, and typeset the underline
    (i.e., caption + legend) to the width of the block (e.g., artwork)
    (if it's a long underline) or centre it (if it's a short one).
(i) This deals with illustrations that are too wide.
(j) This deals with spacing around a block (e.g., artwork)
(k) This deals with spacing above/below the underline (i.e. caption
    + legend).
(l) TWO ILLUSTRATIONS THAT ARE EACH LESS THAN HALF THE TEXT-WIDTH
    ARE TO BE PLACED SIDE-BY-SIDE.
(m) WHEN TWO ILLUSTRATIONS (of "not too dissimilar depth") ARE
    PLACED SIDE-BY-SIDE, THEY ARE TO BE PLACED VERTICALLY SO THAT
    EITHER THE BASES OF THE BLOCKS (E.G. ARTWORK) ARE LEVEL
    OR     THE BASES OF THE UNDERLINE (I.E. CAPTION + LEGEND) ARE LEVEL.
(n) This deals with running heads.
(o) This deals with landscape.
(p) "FOR MULTI-COLUMN MAKE-UP, THE ORDER OF ILLUSTRATIONS IS TO
    BE MAINTAINED WHEREVER POSSIBLE".  I'm not sure what this means.
    It seems to allow both
      Fig 1.  Text
      Fig 2.  Text
       Text   Text
    and
      Fig 1.  Fig 2.
      Text    Text
      Text    Text
(q) This deals with making the underline (i.e. caption plus legend)
    slightly less wide than the block (e.g. artwork), if such adjustment
    is necessary to improve the appearance of the underline.
OTHER GURUS
Leslie Lamport & Margery Cantor
Appended to the notes distributed on the 14th was some mail
from Leslie to Frank about a conversation with a Margery Cantor.
She suggested aligning the tops of figures [whereas Hart suggests
aligning the bottom of the (wood-)block or the bottom of the
underline (e.g. caption + legend)].
Leslie also suggested a few other things, in particular,
that the user will have to arrange for a figure to have
a certain width, perhaps by using a minipage.

Ruari McLean
On page 137 of
"The Thames and Hudson Manual of Typography"
Thames and Hudson, 1980,  ISBN 0-5000-68022-1,
Ruari McLean suggests the following:
If a book consists of a number of text pages ... and the pages
include a number of squared-up illustrations, all different shapes,
the unity of the book is immediately threatened.  The illustrations
should all have the same width as the type area ... If they can
also, with captions, make the same depth as the text, so much the
better ...
If a book is designed with wide foredge margins, and it is
strongly desirable to make an illustration wider than
the type measure, it can be taken into the margin, provided
it is conspicuously wider than the text;  but if it is done
more than once, the larger illustrations should all be of the
same width.
... metal blocks ... flanges of about 1/8 in ...
Half-tone illustrations ... illustration area in harmony with
the type area ...
... undesirable to lay down a set of rules ... But a few "rules
of thumb" are helpful ...
If two pictures are to be placed one above the other, they
should either be identical in width or conspicuously different.
If they are unavoidably of nearly the same width, they should not
be aligned on one side, but placed full out to left and right
respectively, so that the comparison is not obvious.
Hugh Williamson
In chapter 14 of
"Book Design", Yale University Press, 1983, ISBN 0-300-03035-5,
Hugh Williamson suggests the following:
If a designer and illustrator can agree on the dimensions of
pictures before work begins, both the typography of the book
and its illustrations are the more likely to succeed.
...
When ... full-page illustrations vary in depth, a standard
position for the last line of the caption, perhaps aligned
with the last line of text on a full page, ... , even at the
expense of varying distances between caption and picture.
...
When a picture and a long caption are to share the area allocated
to illustration, the dimensions of the picture may depend on
those of the caption, ... way to determine illustration size

will be to have such captions set ... first.
...
A conspicuously rectangular illustration, ... , is usually
placed with one or more of its edges in exact alignment
with other elements of the opening.
...
act of reading favoured by placing the illustration between
two paragraphs ... to divide a word immediately before and
after an illustration is clumsy.
...
Captions often look neatest when they are no wider than
the illustration under which they are to appear.
John Miles
In "Design for Desktop Publishing" (Gordon Fraser, 1987, ISBN 0-86092-097-6),
John Miles gives some examples of grid-based designs, e.g., an A3 design
that has 5 columns each of 12pc with a 1pc gutter between each column
(page 38).  Presumably, such a design will work best if artwork is
supplied 12pc, 25pc, 38pc, 51pc or 64pc wide.
IMPLICATIONS
I don't say that the opinions of Hart, McLean, Williamson, etc. represent
"what LaTeX 3.0 should do".  But their opinions seem to represent
the sort of thing that LaTeX 3.0 should allow a designer to implement.  To
this end, it seems to me desirable that LaTeX 3.0 and the designer should be
working in terms of the same concepts.
From the above, it appears that the following concepts may be fundamental:
In old-fashioned terms "the positions of the edges of the wood/metal block
holding line-art or half-tone" or, in PostScript terms, "the
bounding box for the file of encapsulated PostScript that is being
\special-ed in at this point to define some artwork"
The designer may have a preference about the sizes of illustrations that
are to be allowed.  The artwork may have to be scaled/cropped to be
one of a designer-approved set of sizes (or widths).
The width of the underline (= caption + legend = 2.09 \caption) may
depend on the width of the ("wood or metal") block or EPS
bounding box (e.g. artwork) [or occasionally vice versa].
The underline (= caption + legend = 2.09 \caption) may need to
move (e.g. to the side of the artwork) if the artwork is narrow.
Alternatively (Williamson), a caption might move to be aligned
with the bottom of the last line of text on a full page.
The underline may actually contain: a caption; a legend; details of source.
They may need to be distinguished typographically.

LaTeX may need "to know" the height and width of the artwork (i.e.,
height/width of wood/metal block, or EPS bounding box), so that it
can align "the edges of the block" with other elements of the page.
In particular, the width of the artwork may determine the width of
the caption.  LaTeX may also need to manipulate the underline (i.e., 2.09
\caption) to side of page (Hart) or bottom of page (Williamson).
LaTeX may need to align the {artwork + underline} by using the edges
of the artwork or by using the top/bottom lines of the underline;
hence LaTeX needs explicit access to these positions.  I have the
impression that LaTeX needs to be working in terms of: "the bounding
box for the artwork"; the width of the {caption+legend+source}, the
baselines of the top/bottom lines of the {caption+legend+source}. I have
the impresssion that the height of the \begin{figure} ... \end{figure}
and the width of an enclosed minipage are unlikely to be good enough.
Therefore, it looks to me as though it would be better for the user (guided
by the designer's preferences) to declare the height/width of the artwork
(rather than the height/width of the artwork+caption) and let LaTeX work out
the height/width of the "artwork + caption" (which could conceivably involve
trying having the caption at the side of narrow artwork).  Then LaTeX
will "know" the width of the artwork and can, if appropriate, typeset
the caption to the same width (whereas if width was specified by
a minipage containing the whole lot, it might be wider than the
actual artwork, and the caption might end up too wide).
So I wonder whether some user interface such as the following might
be appropriate:
\begin{floatfig}     %  A golden opportunity to phase "figure" out.
                     %  My users often don't understand that it "floats"
                     %  until they see it happen, then say they don't like
                     %  it!  A more meaningful name would help understanding.
\artwork{144pt}{216pt}        %  Here I assume that the dvi... command
\special{dvi... 144 ... 216 } %  can scale, and that we want to
                              %  take advantage of this feature in this
                              %  case to scale the artwork to fit within
                              %  a 2in x 3in bounding box, perhaps because
                              %  the designer has expressed a preference
                              %  for artwork that is 2in wide.  (If not,
                              %  the user would simply specify
                              %  Xpt and Ypt in the \artwork command
                              %  where X and Y are taken from the
                              %  BoundingBox comment in the EPS file.)
                              %  A \begin{picture}(X,Y) could be used
                              %  instead of \artwork ... \special ...
                              %  if the picture environment is being used
                              %  for a drawing.
\figcaption{...}     %  LaTeX now knows the width of the artwork,
                     %  so can, in principle, try moving the caption

                     %  to the side of narrow artwork or typeset the
                     %  caption to the same width as the artwork.
\end{floatfig}
INCREASING THE PROBABILITY OF SUCCESS
The discussion on the 14th postulated a situation in which it might
be necessary to place arbitrary collections of figures of arbitrary
size.
Obviously "it would be nice" from an algorithmic point-of-view
to have an algorithm than could handle such a situation.  But:
*  In practice, a lot of placement problems may be easier than the
   general problem.  Artwork may actually be a graph off a graph-plotter
   (produced using the plotting package's default plot-size and scaled
   using a constant factor):  all plots will tend to be the same size.
   Photographs will tend to be supplied in standard sizes.
   [On the other hand, chemical artwork may tend to produce
   diagrams of differing sizes, because chemists may want the
   bond-length constant throughout the document, rather than
   the "structure-diagram size" constant.]
*  The designer may express a preference for particular sizes (or at
   least particular widths) for line-art, half-tones, etc.
   A designer might even say that "arbitrary collections of figures
   of arbitrary size" threaten the unity of the book (McLean).
*  dvi-printing commands often allow for the scaling of a \special-ed
   piece of artwork to a particular size.
So a lesser objective than "arbitrary collections of arbitrary sizes"
might deal with the most common cases (and might conceivably lead to
better design).
There would have to be a convention whereby the designer could
pass on his/her preferences about the sizes in which artwork should
be supplied.  I suppose that comments in the style-file would do.
For example, for Miles's A3 design, a comment in a style-file might say
      In this design, LaTeX's algorithm for figure-placement works best
      with figures that have the following widths: 12pc, 25pc, 38pc,
      51pc, 64pc.
In a design for a single-column book, the comment in the style-file
might recommend "magic numbers" chosen as "full-width", "a bit less than
half full-width" etc.  Generally, the user could be told
      If you have a choice, we advise you to use one of the following
      widths for best results with this design ...
Then:
*  users will know how to use dvi-printing commands' abilities to
   re-scale artwork.  (Otherwise, they'd probably just use whatever
   size picture happens to come off their graph-plotter, out of their

camera, etc.)
*  LaTeX is likely to get given artwork of "a recommended width" that
it can make a good job of placing, since much of the testing of
the placement algorithm and style-files would be with artwork of
"recommended widths".

Tables

4.1 Decimals

The recommendations that I’ve seen for table layout ([3, p. 332], [4, p. 164], [7, p. 158], [9, p. 147], [17, sec. 7.2]) mention (or imply) “making sure that, in a column of numbers, the decimal points line up”. Thus “alignment of decimal points” seems a “fundamental requirement”, at least as important as tabular’s current l, c and r, not an “optional extra”.

I realise that the L^{A}TEX 2.09 manual [1] shows a couple of tricks that can be used (zero-width boxes, page 98; high–low, page 182), and that

Frank Mittlebach has the >{\centerdots}c<{\endcenterdots} idea [18, p.

299]. But these techniques seem unnecessarily subtle for such a fundamental requirement.

Couldn’t there be something as easy as l, c and r for “alignment on decimal points”? Bell Labs. managed to make “lining numbers up” easy in tbl [19, p. 147] years ago!

4.2 Rules, i.e. lines

I’ve tried to use L^{A}TEX 2.09 to typeset some of the tables shown as examples by gurus of table layout ([7, p. 157], [17, p. 6]), but have had only partial success.

It looks as though, for success in laying tables out as the gurus recommend, one needs:

• a choice of thickness for \hlines ([7, p˙156], [17, p. 4]). If one follows [17], the choice would be between 0.5pt, 1pt, 2pt and 4pt.

• an option of specifying that adjacent \clines should not join up.

4.3 Notes

The facilities provided by L^{A}TEX 2.09 (“put a tabular inside a minipage and use \footnote”) do not seem particularly suitable for providing notes about tables in the way advocated by the gurus.

• The gurus seem to typeset notes no wider than the table [3, ch. 12], [9, p. 147]. However, if one has to declare the minipage before typesetting the tabular, the minipage must be wider than the expected width of the tabular, so the footnotes will generally be wider than the table.

• The gurus advocate putting “general notes” before “specific notes” ([3, p. 333], [4, p. 164]). There does not seem to be an easy way of getting “general notes” out in the style that \footnote will use for “specific notes”.

• Having to redefine \footnoterule (to give a null rule) is an added irritation.

It might be better if:

• the manual for L^{A}TEX 3.0 refrained from suggesting “put a tabular

inside a minipage, use \footnote and re-define \footnoterule”

• L^{A}TEX 3.0 made the width of the most recently typeset tabular available in a variable, so that the user could declare a minipage or a tabular of the same width (or a related width) immediately below, to hold the notes.

• there was some means whereby the user could tell L^{A}TEX to base it’s decision about tabular-width on the rows that hold the table-proper but could let the tabular continue below the table-proper. Notes could then be provided in “extra rows” that had the same width as the table-proper.

It would be even better if the user specified the table and its surroundings in terms of the logical elements “title”, “subtitle”, “table proper”, “source notes”, “general notes”, “specific notes”, “probabilities” (as suggested in section 3.2.3), and L^{A}TEX 3.0 and the sty file took care of the detailed layout.

4.4 An e-mail addition

From:     David Rhead ...
Subject:  Multi-line cells in tables
Date:     17 Aug 90 12:01:50
I've been browsing through some books that describe, among other things,
how one should typeset tables.  The gurus seem to regard a table entry
that consists of a "hanging paragraph" as fairly standard, either because
they explicitly advocate it, or because they simply do it.  See:
Jan V. White "Graphic Design for the Electronic Age",
      Watson Guptill, 1988, page 147 (the bit about "first line hanging
      indent").
"Chicago Manual of Style", Chicago University Press, 1982,
      pages 338 and 393.
Judith Butcher "Copy-editing", Cambridge University Press, 1981,
      page 136.
"Hart's Rules", Oxford University Press, 1983, pages 34-35 and 158.
(I think that it is desirable that the LaTeX 3.0 defaults should generally
correspond to standard publishing practice.)
Unfortunately, it does not seem to be simply a matter of "In this
\documentstyle, tabular's p{...} has to give hanging paragraphs".
Some respectable books seem to have left-justified paragraphs
in some tables and hanging paragraphs in others.  (E.g. in spite of the
hanging paragraphs mentioned above, Chicago has left-justification on pages
176-177 and Butcher has left-justification on pages 68-69.)  I can't detect
any underlying "logical structure" that corresponds to the difference between
hanging and left-justified within the same book.
In LaTeX 3.0, might it be worth doing one of the following?
Either provide something like p{...} [perhaps h{...} for "hanging"]
      that gives a hanging paragraph rather than a left-justified paragraph
or, if one takes the point-of-view that books that use more than one
      visual representation for the logical element "paragraph in cell"
      are wrong (now matter how respectable and esteemed they are),
      at least put a comment near the code for "paragraph in cell" in
      the standard style files (or wherever) saying how one can change
      things so that all "paragraphs within cells" are hanging.
Perhaps one might be able to achieve the same effect by using
Frank Mittelbach's >{\parindent=...}p (if \parindent=-1em works),
as described in TUGboat, vol. 9, page 299, but it looks as if "hanging
paragraph within table cell" is sufficiently standard among gurus that it
would be worth providing explicit support in some way.
Incidentally, most gurus mentioned above advocate/do "paragraph within cell"
ragged right but hyphenated.  (The exception is Butcher who has ragged right

most of the time but flush right on page 289, possibly because the column on
page 289 is wider than columns elsewhere.)  Therefore, it would seem wise
to make ragged right the LaTeX 3.0 default for "paragraph within cell".
If a cell contains figures which are described by a multi-line table-stub,
how would one get the figures aligned vertically with the bottom line of the
stub?  I have in mind things like
      Infected trees, %   1.63     0.9   20.3
      Chi-square for
        observed values   7.83   11.09   4.98
(on page 158 of Hart's Rules) and
      Interest and discount on other
      lending in sterling              53   38   54   98   141   168   211
(on page 65 of John Miles's "Design for desktop publishing").  Would
one use [a hanging variation] of {b{...}r} or {b{...}t{...}} (where b{...}
and t{...} are as described in Frank Mittelbach's article)?  Again, this
convention for alignment seems sufficiently common that I think it should
be available via a simple syntax if it isn't already covered by existing
proposals.

Lists of references; bibliographies

5.1 Title

The L^{A}TEX 2.09 “standard styles” seem to assume that a “bibliography” is merely a long “list of references”. However, many people use “bibliography” to mean something different from “list of references”. There may also be situations where:

• an author wishes to use something like thebibliography for “Select Bibliography” or “Further Reading” ([3, p. 420], [4, p. 182]).

• a document might have more than one thebibliography-like list: for example (a) “References” then “Further reading” [4, p. 192] (b) “List of references” then “Bibliography” [32, p. 22].

It seems to me that it is important to enable the author to give thebibliography a title that accurately describes its contents.

Consequently, I think that the approach take in german.sty (in the Aston archive) is mistaken. For example, having \def\refname{References} and \def\bibname{Bibliography} for English will tend to perpetuate the mistaken assumption that “an article has a list-of-references while a report or book has a bibliography”, and hence to perpetuate inaccurate titles.

I think that:

• it would be better to have just one default title per language, irrespec-

     tive of \documentstyle

• the default title should be the usual description of a list from which

     one \cites. For English, we might have \def\refname{References}

as the default.

• there should be an easy mechanism (such as that shown in Figure 5.1) to allow the author to replace the default title.

5.2 Schemes supported

In scientific, technical and medical work, the two most important citation schemes are “reference by number” and author-date. In the humanities, references may be given using the “short form” scheme (often in footnotes). All three schemes are supported by the ISO standard. See [30, p. 25] and [28].

The definition of \cite and \bibitem in L^{A}TEX 2.09 supports “reference by number”, but anyone wanting to use a standard author-date scheme:

• has to re-define \cite, to get “(” and “)”

• has to re-define \bibitem, to suppress the [key] printed by default

• will probably want a command that delivers just the “date” part when the author’s name occurs naturally in a sentence. See [28, sec. 9.4].

This generally means that people wanting to use author-date:

• are unable to use standard programs such as lablst which assume the original L^{A}TEX 2.09 definitions

• have to specify “I’m using author-date” in two places: as a style-option

     in their \documentstyle command, and in their \bibliographystyle

command. This seems unreasonable.

Similarly, the “short form” scheme is unsupported.

Can anything be done in L^{A}TEX 3.0 to give author-date and “short form” the same status as “reference by number”? Some ideas will be found in Appendix D.

%  This style-option file, thebib-plus.sty, is based on code taken from
%  article.doc and report.doc/book.doc.
%
%  It differs from the standard code in that:
%  - a command \bibtitle is defined, to generate the bibliography title
%  - a command \bibheadtitle is defined, to generate the text to be used
%    by \pagestyle{headings}
%  - \bibtitle and \bibheadtitle are set to generate "References" and
%    "REFERENCES" respectively.
%  - an \addcontentsline{toc}{...}{\bibtitle} command is issued.
%  Thus:
%  - the default bibliography title is "References", but can be changed by
%    having \renewcommand{\bibtitle}{...} before \begin{thebibliography}
%    or \bibliography
%  - the default page header text is "REFERENCES", but can be changed by
%    \renewcommand{\bibheadtitle}{...}
%  - \tableofcontents will give a table of contents that mentions the
%    bibliography.
%
%  Define default title and page-head text.
%
\def\bibtitle{References}
\def\bibheadtitle{REFERENCES}
\@ifundefined{chapter}{\def\thebibliography#1{\section*{\bibtitle\@mkboth
  {\bibheadtitle}{\bibheadtitle}}
  \addcontentsline{toc}{section}{\bibtitle}\list
  {[\arabic{enumi}]}{\settowidth\labelwidth{[#1]}\leftmargin\labelwidth
    \advance\leftmargin\labelsep
    \usecounter{enumi}}
    \def\newblock{\hskip .11em plus .33em minus .07em}
    \sloppy\clubpenalty4000\widowpenalty4000
    \sfcode`\.=1000\relax}}%
{\def\thebibliography#1{\chapter*{\bibtitle\@mkboth
  {\bibheadtitle}{\bibheadtitle}}
  \addcontentsline{toc}{chapter}{\bibtitle}\list
  {[\arabic{enumi}]}{\settowidth\labelwidth{[#1]}\leftmargin\labelwidth
    \advance\leftmargin\labelsep
    \usecounter{enumi}}
    \def\newblock{\hskip .11em plus .33em minus .07em}
    \sloppy\clubpenalty4000\widowpenalty4000
    \sfcode`\.=1000\relax}}

Figure 5.1: A way to allow easy changes to the “list of references” title

5.3 Consequences for BibTEX

The situation mentioned in section 5.1, where a document may have more than one thebibliography may have particular implications for BibTEX. A document might have “References” to list all the works cited, and “Further reading” to gives further details of some of them.

In L^{A}TEX 2.09 terms, one might need 2 \bibliographys with 2 different \bibliographystyles. Does one need the “bibliography style” to be an argument of the \bibliography command, so as to allow for this possibility?

5.4 Subdivisions

Except when using an author-date scheme, an author might reasonably wish to subdivide a reference list ([3, p. 425], [4, p. 183]). Can some mechanism (\subbibliography{...}?) be provided so that the author can place a subheading in a reference list to indicate the start of a new category? (It would be nice if the mechanism allowed a bst file programmer to sort by “category field”, and then write an appropriate category heading to a bbl file, so that BibTEX users could potentially have subdivided reference lists too.)

5.5 Note

See also Appendix D.

“Over-long” headings

Although the L^{A}TEX 2.09 manual [1, p. 126] explains that giving an overlong heading argument in sectioning commands, \caption, etc. can cause a “buffer size” error, it seems perverse that an error can occur (for example) because of the size of a “list of tables” string when the user hasn’t actually given a \listoftables command.

Could L^{A}TEX 3.0 deal with over-long arguments in sectioning, \caption, etc., commands as follows?

• Instead of regarding heading as the default toc-entry, text to go in page-header, lot-entry, lof-entry, etc. [1, p. 158], regard just the first 50-or-so characters of heading as the default toc-entry, etc.

• If the user supplies a heading that is more than 50-or-so characters:

– If there is no [lst-entry] argument, put the first 50-or-so characters of heading to the aux file, followed (if truncation was necessary) by \dots or “etc.” or some other truncation indicator. The chances are that the user doesn’t want an entry in the tableof-contents, list-of-tables, list-of-figures, or hasn’t thought about the implications of the length of heading for the table-of-contents, etc. The user may well be mis-using \caption for (in a table) the title and notes or (in a figure) the caption and legend. The L^{A}TEX 2.09 manual ids right when it says [1, p. 126] that “such a long entry won’t help the reader”. The first 50-or-so characters will give the user something that “will do” until they’ve thought about what they really want.

– If truncation was necessary, send a message to the user’s terminal and log file saying, for example, “Warning: Your caption is

more than 50 characters long.  Use \caption's optional
lot-entry argument if you intend to use \listoftables”.

– Allow the user to supply a lst-entry of up to the 2.09 limits if they do it explicitly through the optional [lst-entry] argument. (They’re explicitly using lst-entry, so they probably know what they’re doing and really want these characters in the table-of-contents or whatever.)

This approach should:

• eliminate the situations where L^{A}TEX gives an error message because of an over-long heading even when heading isn’t actually going to be printed anywhere where its length will be a problem

• give the user an authoring problem “How can I improve this stopgap list-of-tables entry that L^{A}TEX has kindly supplied?” rather than a debugging problem “Why is user-hostile L^{A}TEX giving me this obscure error message?”.

Page headers

Suppose that, in a book:

• the design requires chapters to start on a right-hand page

• chapter i ends on a right-hand page, so L^{A}TEX must produce an empty left-hand page before starting chapter i + 1.

In this situation, publishers seem to assume that the empty left-hand page is neither in chapter i nor in chapter i + 1 (see, for example, [3, 1]) and hence that the empty page should have no “Chapter … ” page-heading. L^{A}TEX 2.09, on the other hand, gives it a page-heading to say that it it is part of chapter i (which looks silly).

Could the L^{A}TEX 3.0 version of book (and similar styles) use \thispagestyle{plain} or \thispagestyle{empty} for a left-hand blank page that separates 2 chapters?

Sections

The definition of \@startsection in latex.tex allows for “run-in headings”.

However, a design often involves some punctuation between a run-in heading and the following text, e.g. “a full stop unless the heading naturally ends in a question mark or exclamation mark”. (See, for example, [3, p. 222–223] and [4, p. 240].) Would it be possible, at L^{A}TEX 3.0, to define the successor to \@startsection in such a way that, if there is to be a run-in heading, the sty file can specify default punctuation between the heading and the following text?

E-mail comments

From:     Michael Downes <MJD@COM.AMS.MATH>
Date:     Mon, 8 Apr 91 16:50:30 CET
A punctuation question:
If a documentstyle design calls for punctuation such as period or colon
after a run-in section heading or theorem-environment heading, the
punctuation can be provided fairly easily [although maybe not without
changing \@sect]. But what do you do in particular cases where something
else (like a ?) needs to be substituted for the default punctuation?
Has anyone else faced this question and found a way to do it that
doesn't require maintaining an array of macros \sectionpunct,
\subsectionpunct, \subsectionpunct, ... to be called in \@sect
using \csname?  (And then you might need conceivably also need
\theorempunct, \corollarypunct, \lemmapunct, \examplepunct, for each
environment created with \newtheorem, or at least for each

theoremstyle---AMS journals typically have four or five different
`theorem'styles.)
My first attempt at a solution has involved adding a macro \endpunct at
the appropriate place in \@sect and \@begintheorem. Then \endpunct can
be redefined in a document when necessary. But there are complications
in scoping and ease of use; for example, this fouls up the nobreak and
afterindent mechanisms:
  {\renewcommand{\endpunct}{?}
   \section{Why abandon the old technology}
   \label{s:who}
   \showthe\everypar
  }
   \showthe\everypar % lost the \@afterheading stuff
   Text begins here.
and if we move the closing brace further to avoid losing \everypar, what
rationale do we give the user?
  ...
  T}ext begins here. % try explaining this!
Then there is the potential for situations like this:
  {\renewcommand{\endpunct}{?}
   \section{Why abandon the old technology}
   \label{s:who}
  \begin{itemize}
  \item Text begins here.
  ...
where an environment occurs at the beginning of a section. The closing
brace can't be put inside the environment without causing an error.
In LaTeX a * option would be the natural way of taking off the ending
punctuation but for sectioning commands the * option is already being
used for something else.
Any suggestions?
Here's another idea:
   \section{Why abandon the old technology\substitutepunct{?}}
with

  \def\substitutepunct#1{\def\protect\endpunct{#1}}
I haven't tested this yet so I don't know whether there would be table
of contents complications in this case.
P.S. In the AMSART documentstyle distributed in the AMS-LaTeX package,
where \subsection and lower level headings are run-in with ending
punctuation, I ignored this issue, thinking that the need to substitute
would be so rare that by the time we needed to face it, a good solution
might have turned up.  Naturally the first or second paper submitted to
the AMS using the AMSART documentstyle had a couple of section headings
ending in question marks!

—

From:     Don Hosek <DHOSEK@EDU.CLAREMONT.HMCVAX>
Date:     Tue, 9 Apr 91, 03:53:00 PDT
Michael brought up the issue of section headings where a period
should be added to the end only if there is not already a
punctuation mark at the end of the given section head. To which I
offer a couple thoughts in order of increasing attractiveness:
 (1) pull items off the horizontal list and decompose them to see
 if there's a punctuation mark at the end
 (2) just before reading the argument containing the section
 title, make !, ? and . active and have them set a flag in
 addition to printing the appropriate mark. If that flag is true,
 don't add a period onto the end yourself.
 (3) take a look at the spacefactor. If it is 3000, we've just
 had end-of-sentence punctuation. It's not too difficult to
 justify the occasional \@ which would be necessary under these
 circumstances.

—

From:     MITTELBACH FRANK <PZF5HZ@EARN.RUIPC1E>
Date:     Tue, 9 Apr 91 18:28:28 CET
Don send some interesting suggestions how to handle
punctuation characters in `certain' situations (not only
headings). But I think only the last might work and even
this one will work only if we define things carefully.

>
>  (1) pull items off the horizontal list and decompose them to see
>  if there's a punctuation mark at the end
Splitting something up at the right end of a token list is
difficult and will not work if the list contains braces and the
like. But if anybody has a good algorithm for this general problem
I'm interested to see it.
>
>  (2) just before reading the argument containing the section
>  title, make !, ? and . active and have them set a flag in
>  addition to printing the appropriate mark. If that flag is true,
>  don't add a period onto the end yourself.
Active characters won't do due to the problems that TeX fixes
catcode whenever it reads arguments. The only possibiltiy I see
when such a strategy is used is to have all these chars *always*
active with a default definition that produces the non active
equiv. But this will slow down processing very much.
Anyway, something like this will be implmented to provide short
references but I think it isn't justified to use this for
punctuation in general.
>
>  (3) take a look at the spacefactor. If it is 3000, we've just
>  had end-of-sentence punctuation. It's not too difficult to
>  justify the occasional \@ which would be necessary under these
>  circumstances.
That's an interesting idea but assumes that punctuation chars
always have a \spacefactor that distinguishs them from other
characters. This means that we need to change \frenchspacing and
I'm not sure what happens then, how many applications depend
on things like this. But it looks certainly promising.

—

From:     Don Hosek <DHOSEK@EDU.CLAREMONT.HMCVAX>
Date:     Wed, 10 Apr 91 00:01:00 PDT
->  (2) just before reading the argument containing the section
->  title, make !, ? and . active and have them set a flag in
->  addition to printing the appropriate mark. If that flag is true,
->  don't add a period onto the end yourself.
-Active characters won't do due to the problems that TeX fixes
-catcode whenever it reads arguments. The only possibiltiy I see

-when such a strategy is used is to have all these chars *always*
-active with a default definition that produces the non active
-equiv. But this will slow down processing very much.
-Anyway, something like this will be implmented to provide short
-references but I think it isn't justified to use this for
-punctuation in general.
... Consider the following simple version of the above algorithim:
\def\section{\catcode`\.=\active\xsection}
\def\xsection#1{\centerline{#1\ifperiod.\fi}\periodtrue\catcode`\.=11\relax}
{\catcode`\.=\active
 \gdef.{\periodfalse}}
\newif\ifperiod
\periodtrue
(I originally wrote this sort of code to handle the dropped E
problem of tex.web). Incidentally, the above code won't quite
work since a period in the middle of the title will gleefully
supress the final period. Perhaps an \ifnextchar sort of setting
would work better.
->  (3) take a look at the spacefactor. If it is 3000, we've just
->  had end-of-sentence punctuation. It's not too difficult to
->  justify the occasional \@ which would be necessary under these
->  circumstances.
-That's an interesting idea but assumes that punctuation chars
-always have a \spacefactor that distinguishs them from other
-characters. This means that we need to change \frenchspacing and
-I'm not sure what happens then, how many applications depend
-on things like this. But it looks certainly promising.
All we need to do is make sure that the chapter titles are set
with \nonfrenchspacing. Simple enough. In any event,
\frenchspacing is really not a command that should appear in a
TeX document anyway (if they need frenchspacing, chances are
they'll be using a special document style anyway).

—

From:     MITTELBACH FRANK <PZF5HZ@EARN.RUIPC1E>
Date:     Thu, 11 Apr 91 15:37:32 CET
>
> ... Consider the following simple version of the above algorithim:
>
> \def\section{\catcode`\.=\active\xsection}
> \def\xsection#1{\centerline{#1\ifperiod.\fi}\periodtrue\catcode`\.=11\relax}

> {\catcode`\.=\active
>  \gdef.{\periodfalse}}
> \newif\ifperiod
> \periodtrue
... all these kinds of
solutions are for the user that understands TeX fine but won't
work in all situations. Consider
  \newcommand{\mysection} 1 {\section{#1} \prefacetext}
and you've lost already. Not that is a good example but you will
find them if you look for them.
> All we need to do is make sure that the chapter titles are set
> with \nonfrenchspacing. Simple enough. In any event,
> \frenchspacing is really not a command that should appear in a
> TeX document anyway (if they need frenchspacing, chances are
> they'll be using a special document style anyway).
Yes, this would work I think it it seems necessary to set the
spacefactor explicitly. What I'm after are general solutions not
these kind of special hacks that solve this problem but create
twenty new ones if other conventions to be implemented.

Lists

9.1 Within paragraphs

In my opinion, someone reading a document stands a better chance of absorbing the contents if they can follow the author’s thought-groupings. As part of this, they need to be able to be clear about where one paragraph ends and the next begins. Hence, I think they need to be able to distinguish between:

1. “list at end of paragraph”

2. “list within paragraph”.

Otherwise, they won’t be certain about whether there is, or isn’t, a paragraphdivision.

With L^{A}TEX 2.09, this is difficult in designs that have un-indented paragraphs, because “the vertical space following the environment is the same as the one preceding it” [1, p. 114]. Case 1 looks identical to case 2.

Could L^{A}TEX 3.0 implement “the text following a displayed paragraph environment begins a new paragraph if there is a blank line after the \end command” [1, p. 165] for both \parindent and \parskip, so that case 1 looks different from case 2 even when \parindent is zero?^{1}

1I’m probably just agreeing here that L^{A}TEX 2.09 has “an actual conceptual bug” [2, p.

684].

9.2 An interrupted enumerate?

Would commands such as \suspend{enumerate} and \resume{enumerate}

be possible, so that the author can introduce some items in a list with one sentence and then introduce further items in the same list with another sentence? It would be easier than getting involved with \setcounter{enumi}.

verbatim

10.1 Vertical space

For lists in designs that have un-indented paragraphs, I generally set \parsep and \itemsep to zero and \topsep to -\parskip, to eliminate all vertical spacing.

The objective is to make it clear where paragraphs begin and end. I think that:

• extra “blank lines” in lists, and around verbatim, will make paragraphdivisions less clear

• the bullet or “label” is a clear enough indicator of where a list-item starts, and the font-change is a clear enough indication of where verbatim starts.

So far, with L^{A}TEX 2.09, I’ve been unable to produce style files that close lists up satisfactorily when a list contains some verbatim. I guessed that, since verbatim uses \trivlist, changing the list spacing should eliminate the vertical space around verbatim too. Unfortunately, I end up with the wrong spacing around verbatim.

I would appreciate anything that could be done in L^{A}TEX 3.0 to make it easier to eliminate the vertical spacing around verbatim, as can be done for lists.

10.2 Horizontal space

It is usual to show computer input/output in verbatim. One might reasonably say that “computer input/output is a type of quote”, and so want to put a particular horizontal space onto each line before verbatim material is printed. In L^{A}TEX 2.09, this does not seem easy (and, because verbatim can’t be used in the definitions of other environments, one can’t define another environment that does it either).

Could something be provided in L^{A}TEX 3.0 that does for verbatim what \leftmargin does for lists?

Borrowing some margin

Problems often arise in publishing practice because certain features (such as tables, figures and equations) are a little too wide for the normal portrait text-width (see [3, p. 343] and [4, p. 161]). A recognised solution seems to be to temporarily increase the text-width [3, p. 343].

Might it be worth defining a standard environment (borrowmargin ?) in L^{A}TEX 3.0 that does this? Figure 11.1 shows a prototype definition. (It would be better to avoid the minipage in this prototype, since it changes footnote numbering and affects the vertical spacing around the enclosed feature.)

%  The borrowmargin environment is intended for dealing with wide objects.
%  It makes an extra 15mm available on either side of the normal text-area.
%  Marginpars are not allowed.
\newdimen\fullwidth
\fullwidth=\textwidth
\advance\fullwidth by 30mm
\newenvironment{borrowmargin}{%
      \def\marginpar{\typeout{Marginpar is not allowed within borrowmargin.}}
      \savebox{\widethingstore}\bgroup\begin{minipage}{\fullwidth}}{%
%
      \end{minipage}\egroup
      \begin{flushleft}
      \hsize=\fullwidth
      \mbox{}\kern-15mm\usebox{\widethingstore}
      \vspace{1mm}
      \end{flushleft}
}

Figure 11.1: A prototype borrowmargin environment

Appendix A

Some e-mail comments about structure

From:     MITTELBACH FRANK <PZF5HZ@EARN.DRUEDS2>
Date:     Fri, 5 Oct 90 17:43:46 CET
I would like to foward a mail from Nico which was a reaction
to some discussions at the Cork meeting.
------------------------ forwarded mail -------------
...
> Front matter
> ------------
> Currently, an article starts with \title, \author and \date
> instructions. In the Elsevier styles I have added \address,
> \received, \revised and \accepted commands, and also a keyword
> environment, similar to the abstract environment. By doing so, we can
> automatically convert a LaTeX-coded document to an SGML-coded
> document.

—

From:     Don Hosek <DHOSEK@EDU.CLAREMONT.HMCVAX>
Date:     Mon, 18 Feb 91 16:36:00 PST
A few items which have been brought up previously, but I'm
willing to bring them up again now that I have some direct
experience with the situation in a production situation:
\chapter*, to make sense, should work as follows:

- A TOC entry should be generated
- Headers should be appropriately modified
+ All sections contained in it should be unnumbered. Not sure
  about "numbered-within" items like equations and figures. The
  situation has yet to come up
Other sectioning commands should work similarly
Items like \tableofcontents and \thebibliography which currently
use \something* to generate heads can use the low-level commands
to get at the header alone. Things would make more sense in the
style files this way as well.
My use of \chapter* is to mark the prefatory sections of a
document. The first non-starred chapter command switches to
arabic numerals for pages.
...

—

From:     Sebastian Rahtz <S.P.Q.Rahtz@UK.AC.SOTON.ECS>
Date:     Tue, 19 Feb 91 13:54:39 gmt
 > \chapter*, to make sense, should work as follows:
 > - A TOC entry should be generated
no, not necessarily. that should be clearly up to the style designer.
I dont want Preface and Acknowledgments in my TOC, thanks, in style B
 > + All sections contained in it should be unnumbered. Not sure
 >   about "numbered-within" items like equations and figures. The
 >   situation has yet to come up
again, this is not a universal rule. it might make sense, though, to
copy the (S)GML world, and have a higher level constrcut like
\begin{prefacesection} and \end{prefacesection} in which, by default,
sections are all * type
isn't Don's requirement catered for by the values of tocdepth and
secnumdepth, used judiciously? one can reset them on the fly, yes?

—

From:     Don Hosek <DHOSEK@EDU.CLAREMONT.HMCVAX>
Date:     Wed, 20 Feb 91 19:14:00 PST
> > \chapter*, to make sense, should work as follows:
>4 > - A TOC entry should be generated

>no, not necessarily. that should be clearly up to the style designer.
>I dont want Preface and Acknowledgments in my TOC, thanks, in style B
Well you might not, but picking five books at random from my
bookshelf (it's so nice having a computer at home, finally),
reveals:
Norman Sherry, _The Life Of Graham Greene, Vol. I_: Various *-forms
in TOC.
Jan White, _Graphic Design for the Electronic Age_: ditto.
Floyd L. Moreland & Rita M. Fleisher, _Latin: An Intensive
Course_: ditto
Leslie Lamport, _LaTeX: A Document Preparation System_: ditto.
Bentley Layton, _The Gnostic Scriptures_: ditto.
I haven't tried, but I suspect that there are no examples of the
opposite practice in my library.
> > + All sections contained in it should be unnumbered. Not sure
> >   about "numbered-within" items like equations and figures. The
> >   situation has yet to come up
>again, this is not a universal rule. it might make sense, though, to
>copy the (S)GML world, and have a higher level constrcut like
>\begin{prefacesection} and \end{prefacesection} in which, by default,
>sections are all * type
>isn't Don's requirement catered for by the values of tocdepth and
>secnumdepth, used judiciously? one can reset them on the fly, yes?
Kind of. You get one tocdepth per document, the one in affect
when \tableofcontents is executed. One _could_exert some effort
in writing \setcounter{tocdepth} commands to the toc, but I'm not
inclined to bother. Setting secnumdepth to a negative number for
the prefatory material, though is a very appealing idea and
solves the problem quite nicely. I think I'm going to change a
few sections of my LaTeX book once I finish this note.
Incidentally, in keeping with established LaTeX usage, I would
suggest that we have single commands to indicate the beginning of
prefatory material and the beginning of the body a la \appendix
and \ps (think about that one... the analagy seems to be the best
way for my students to understand what \ps is about when I teach
the LaTeX course).

—

From:     David Rhead ...
Date:     20 Feb 91 16:20:30
BACKGROUND
Most users of TeX and LaTeX 2.09 are not committed TeXies: they just want
typesetting that looks as if it was produced by traditional means.  If
TeX/LaTeX are working in terms of structures/concepts/practices that are
different from those adopted in mainstream publishing, authors (and support
staff) are going to be forever translating from LaTeX-speak to publisher-speak.
I appreciate that TeX and LaTeX 2.09 were produced in limited time, and that
if their authors had had to completely research structures/concepts/practices
before they started coding, they might never have had time to write any code.
But if, with the move to LaTeX 3.0, there are opportunities to phase out
the old ways of doing things and to phase in new ways of doing things,
I think that it is worth taking these opportunities to move LaTeX in the
direction of traditional mainstream publishing.  Then the communications
problems between designers and computer-people will be reduced, and authors
will be able to get what designers intend fairly painlessly.
Thus, I tend to think that, if some aspect of LaTeX 2.09 is out of line
with traditional mainstream publishing practice, then it is LaTeX 2.09
that is in "the wild blue yonder".  Admittedly, it's not always easy to
identify what is fundamental in traditional mainstream publishing, but its
better TO TRY than to risk leaving LaTeX out on a limb.
\CHAPTER*
As far as I can see, there is no concept in traditional mainstream
publishing practice that really corresponds to \chapter*.
The relevant concepts seem to be "front matter" (or "preliminary pages",
or "prelims") and "back matter" (or "end pages" or "back matter").
Within these divisions, there are top-level units (Preface,
Acknowledgements, Glossary, References, Index), but they aren't
really chapters.  See, for example,
"Chicago Manual of Style", Chicago University Press, 1982,
      pages 4-5;
Judith Butcher, "Copy-editing", Cambridge University Press, 1981,
      chapters 7, 8, 9;
Hugh Williamson, "Methods of Book Design", Yale University Press,
      1983, chapter 8;
Ruari McLean, "The Thames and Hudson Manual of Typography",
      Thames and Hudson, 1980, chapter 10;
Jan V. White, "Graphic Design for the Electronic Age", Watson-Guptill,

      1988, pages 156-161;
John Miles, "Design for Desktop Publishing" Gordon Fraser, 1987,
      pages 58-61.
This divergence between LaTeX 2.09 and traditional publishing practice
seems to be what causes: Don's problems; Sebastian to have to
countenance resetting tocdepth and secnumdepth "on the fly" (OK for
Sebastian, but I think that it would just put ordinary users off);
ordinary mortals to mess around with \pagenumbering{roman} and
"\pagenumbering{arabic} right after the first \chapter command".
Don wants to signal "prefatory sections" (i.e., what a publisher would
call "front matter", "preliminary pages" or "prelims").  In his design,
"front matter" has:  TOC entry; modified headings; page numbering
in roman; equations and figures unlikely.  I presume that modifying the
effect of \chapter* to meet Don's requirements would violate the
decision that "LaTeX 3.0 should be able to process LaTeX 2.09 input
files", so we'd be talking in terms of phasing something else in to
meet his requirements.
Sebastian's design happens to treat "front matter" differently from Don's,
but I think he still has a chunk of document that would be recognised
as "front matter", etc.
I agree with Sebastian that "a higher level construct" would be appropriate.
I'd suggest something like \begin{prelimpages} ... \end{prelimpages}
\begin{maintext} ... \end{maintext} \begin{endpages} ... \end{endpages}
[If you want a precedent, I think that MIT-press-book.sty envisages that
the author will go \begin{frontmatter} ... \end{frontmatter}.]
Then the user wouldn't have to bother about: the \chapter/\chapter*
distinction; \pagenumbering; \addcontentsline.  Such details should
be taken care of by the designer, via the style file.  The designer
will be thinking in terms of "preliminary pages" and "end pages",
the user will "tell" LaTeX (via the environments) which pages
are "preliminary pages" and "end pages", the user will have to think
in the terms in which his/her publisher is thinking but won't have
to bother with how to coerce LaTeX to make the "preliminary pages"
and "end pages" conform to the publisher's design.  As a bonus, "preliminary
pages" could have a \countN set appropriately to help page selection, e.g.
\count2 = 1, 2, 3 might mean "preliminary pages", "main text", "end pages".
...

—

From:     N.POPPELIER@NL.ELSEVIER
Date:     Thu, 21 Feb 91 09:03:38 +0000

I agree with David Rhead that there is a definite need for a division of a
book in front matter, body and back matter. In fact, last summer I suggested
that we start defining the structure of every class of documents for which we
will make a document style in SGML-like terms. Existing SGML dtd's for books
already contain the division David suggested.
...

—

From:     CA_ROWLEY@UK.AC.OPEN.ACS.VAX
Date:     Mon, 8 Apr 91 13:22:53 GMT
This message is partly ... and partly to
say that David Rhead's suggestions seem to be a very reasonable
outline programme for when we get to that stage.
However, in the last sentence, he says:
> I'd suggest a move towards the "publishing-industry standard" structures.
>
The only such standard I know of at present for the logical structure
of books is the AAP's DTD.
I should therefore like to hear people's views as to whether we should
be thinking about doing this and, also, from those more familiar with
it (Nico!), whether it would be feasible/sensible etc.

—

From:     Michael Downes <MJD@COM.AMS.MATH>
Date:     Mon, 8 Apr 91 18:46:30 CET
I am currently working on a documentstyle with a slightly weird design.
It's for a book consisting of chapters by various authors; the chapters
are collected from various sources, and may be previously published, or
given as a talk at some conference but not previously published, or
written specifically for the current book. I am wondering about
questions such as the following.
---Structure of a chapter: Make it exactly like a journal article?
  \title{Chapter title}
  \author{Chapter author}
  \affiliation{Author's university or institution}
  \maketitle

Or use \chapter instead of \title?  Each chapter is numbered with a big
chapter number, which wouldn't normally be the case in, say, a
conference proceedings volume.
Maybe it's only a question of semantics; but if the publisher wants to
call them chapters, my first choice is to use the \chapter command, to
avoid confusing end users and the publisher's staff.
However, this immediately raises another difficulty: \chapter
does not normally have an associated \maketitle command, but the format
of the author names and affiliations is such that typesetting
each one as it comes along, and getting it in the right place,
would be technically difficult. So I am leaning to the format
  \chapter{Chapter title}
  \author{Chapter author}
  \affiliation{Author's university or institution}
  \makechaptertitle
But this is a little out of synch with the LaTeX manual.
---Processing each article separately or processing them together using
\include. The main benefits of processing them together seem to be (a)
less chance of error in the page numbering and (b) cross-references
between chapters---but this is unlikely to be applicable for
independent articles.  The main drawback of processing articles by
different authors together is the possibility of a global change in one
article affecting subsequent articles in undesirable ways. Local
changes can of course be limited to the original article by enclosing
it in a group. In AMS journals we have experimented with processing
all the articles of an issue together, and encountered undesirable
global changes, as well as save stack overflow from the extra
level of grouping, often enough that eventually we went to separate
processing, with the sequence of articles controlled in a VAX/VMS DCL
command procedure, and with the page number being passed from one
article to the next using \write and \read.
---There is also a possibility in any given volume that certain commands
might be required both within a chapter and at the outer level:
\tableofcontents, \bibliography, \appendix. Something of course
can be managed, but does anyone have previous experience and/or
suggestions on how the user interface ought to look? Should I
provide \tableofcontents and \chaptertableofcontents,
\bibliography and \chapterbibliography, etc.? Or how about
putting commands such as \frontmatter, \middlematter, \backmatter
in the driver file that would change the effect of \tableofcontents,
\bibliography, and \appendix commands? I.e.,

\documentstyle{X}
\begin{document}
% front matter beginning here
\tableofcontents
\include{preface} % containing \chapter{Preface}
\include{notation} % containing \chapter{List of Notation}
\middlematter
\include{chapter1} % normal chapters, maybe containing
\include{chapter2} % their own \tableofcontents, \appendix,
\include{chapter3} % or \bibliography commands.
...
\backmatter
\appendix % switch to appendix format
\include{app1} % \chapter{Appendix title}
\include{app2}
\include{biblio} % \bibliography
\include{index} % \begin{theindex}...\end{theindex}
\end{document}

---On another tangent, the way the \appendix command works has always
seemed slightly odd to me. It would seem more natural to make \appendix
a direct substitute for \section (in an article documentstyle) or
\chapter (in a book documentstyle). Maybe in LaTeX 3.0, with its
attribute handling, we'll be able to say something like
\section[variant=appendix]{...}
?
The chief difficulty seems to be resetting the counter to 1, for the
first such command, and incrementing it naturally thereafter. Other
changes would be straightforward.

—

From:     bbeeton <BNB@COM.AMS.MATH>
Date:     Mon, 8 Apr 91 19:42:51 CET
mike downes has suggested the term \middlematter .  the term
\bodymatter is generally accepted among the sgml people i hang
around in standards meetings, so i'd like to suggest that as a

substitute.

—

From:     N.POPPELIER@NL.ELSEVIER
Date:     Fri, 12 Apr 91 15:10:43 +0000
Chris more or less invited me to comment on the point of publishing-industry
standards and SGML. By coincidence, I have been looking at these things the
past week; next week I will be in CERN working on the AAP dtd and conversions
SGML <--> LaTeX.
O.K., just a quick reply. Sorry haven't got more time. You can skip
everything of this message if you wish, except the part with the ! in
column 1.
There are no real publishing-industry standards: there is a group of
publishers who support the AAP dtd, but the AAP dtd's for scientific articles
and books are not always suited to the needs of publisher X. Therefore,
publisher X adapts the AAP dtd to his purposes and (hopefully) tries to stay
as close to the original as possible.
[The original, by the way, contains several errors, as was pointed out in
EPSIG News by Derek Coleman and Jan Bleeker of our company.]
! Nevertheless, the AAP dtd's for scientific articles and books serve as good
! starting points and I would suggest that we try to create document styles
! that are as close to these dtd's as possible.
What you will immediately notice when you try this, is that they contain much
more elements and sub-elements, especially in the front matter. The
front-matter elements of LaTeX's standard document styles are totally
insufficient. I can understand Michael Downes' attempts at refining this
structure -- in fact I've been doing the same here.
[Another interesting initiative in the SGML field is the TEI initiative,
which is an initiative from the area of humanities and linguistics. A
different approach, but draft version 1.1 of TEI P1 contains interesting
thoughts.]

—

From      David Rhead ...
Date:     12 Apr 91 20:46:08
Here are a few comments about Chris's and Nico's postings.

Martin Bryan gives a DTD for textbooks in Appendix C of his "Author's
Guide to SGML".  I also seem to remember that the British Library had
a project that was defining some DTDs (but I could have remembered wrongly).
Irrespective of whether there are formal standards yet (in terms of SGML
DTDs, etc.), there does seem to be an informal consensus about "the
structure of a book" within the publishing industry.  See, for example,
      Hugh Williamson.  "Methods of book design", chapters 7 and 8.
      Judith Butcher.  "Copy-editing", chapters 7-10.
      Ruari McLean.  "Thames and Hudson manual of typography", chapter 10.
      Jan V. White. "Graphic design for the electronic age", pages 157-161.
      "Chicago Manual of Style", pages 4 and 5.
With a bit of luck, we'll find that the emerging DTDs are consistent
with the previous informal consensus.  For example, if the AAP and Chicago
University Press both manage to produce books in the end, they must both
be talking about the same thing!
As regards the "more elements and sub-elements, especially in the front
matter", I think that this is a price worth paying if it means that
LaTeX and the .sty file can relieve an author of more of the work.
With more accurate knowledge of the structure, LaTeX could take care of
details like:
- when to have page-numbers, and whether they should be arabic or roman
  (rather than authors having to mess about with \chapter*,
  \addcontentsline, \pagenumbering{roman}, \pagenumbering{arabic}
  "right after the first \chapter command")
- when to have table-of-contents entries.
This would help "free authors from formatting concerns to allow them to
concentrate on writing".

—

From:     CA_ROWLEY@UK.AC.OPEN.ACS.VAX
Date:     Sat, 13 Apr 91 16:01:11 GMT
Nico writes (apart from the very useful earlier parts of his message):
> [Another interesting initiative in the SGML field is the TEI initiative,
> which is an initiative from the area of humanities and linguistics. A
> different approach, but draft version 1.1 of TEI P1 contains interesting
> thoughts.]
[TEI = Text Encoding Initiative]
I have not read this document but knew of its existence:
my reason for not mentioning it in my message is that my understanding
of the TEI project is that it is intended to code a far larger range
of "texts" (in particular, historical material) and a far larger range

of properties of those texts (eg what prining press and which
compositor produced them) than any system aimed at typesetting
documents would need to encompass.
Nevertheless, I am sure it contains some matter of interest.  Is it
any easier to study than the AAPs DTD?

—

From:     David_Rhead@UK.AC.NOTTINGHAM.CCC.VME
Date:     Sun, 14 Apr 91 17:36:30 BST
Barbara suggested \bodymatter (as apparently used by SGML people)
as an alternative to the \middlematter used in one of Michael's examples.
The publishing gurus I mentioned in my last contribution use the terms:
"The text" or "main text" (Williamson, p. 180)
"The text" or "main book" (Butcher, pages 119 and 130)
"Main text" (McLean, p. 157)
"Text" (White, p. 159)
"The text" or "text proper" (Chicago, pages 4 and 22)
Would \maintext be worth considering as a command-name?  (If we use existing
publishing-industry jargon rather than defining new jargon, the end-user
gets spared the job of interpreting new jargon in terms of old jargon.
Unless, of course, the publishing industry's jargon is in the process of
changing, in which case we'd be better off using what they are changing to.)

—

From:     David_Rhead@UK.AC.NOTTINGHAM.CCC.VME
Date:     Mon, 15 Apr 91 09:54:21 BST
Michael raised the question of a book that consists of contributions by
various authors.
Obviously, at the moment, he has to do whatever is necessary to get this
particular book published.  However, it would be nice if, in the long term,
there was a recognised way of producing these things.  So I'll comment on
the questions he raised from a long-term point of view, even if the
comments are not relevant to his short-term difficulty.  It seems part
of the question of a "conference-proceedings style" that I think Frank
mentioned informally at Cork (unless I've remembered wrongly).
Generally, I think that having jargon/structures designed for one situation
and then mis-using them in another situation will tend to give confusion
and trouble.  If computer-people and publishing-people use different jargon,

everyone is going to waste time translating from one to the other.  If
they use the same jargon (in this case "chapter") but mean different things,
that's even worse!
In this situation, "being out of synch with the LaTeX manual" is only
to be expected, since the LaTeX 2.09 standard styles weren't really
designed to handle multi-author works.  Using \author when its not
clear whether the author is the "overall editor" or the "author of this unit",
or \chapter when the unit concerned isn't what everyone thinks of as a chapter,
seems like asking for trouble.  The fundamental problem seems to be that
the 2.09 standard styles don't include one that is designed for dealing
with multi-author works.
So, what's required?  Judging by the Chicago Manual of Style and Butcher's
"Copy-editing", we are talking about "multiauthor works", that consist
of "contributions" which are written by "contributors", where the whole
thing is made into a book-like object by a "volume editor".  So an
appropriate set of jargon/structure might go like:
      \documentstyle[...]{multiauthor}
      \volumeeditor{...}
      \begin{frontmatter}
      ...
      \end{frontmatter}
      \begin{maintext}
      \begin{contribution}
          \contributor{...}
          ...
      \end{contribution}
      \begin{contribution}
          \contributor{...}
          ...
      \end{contribution}
      \begin{backmatter}
      ...
      \end{backmatter}
"Contribution" sounds like a word that is already in use in this context, so
that people in the same situation as Michael and his publishers could use it
to avoid confusion about "the publishers calling the units chapters, but
\chapter not being appropriate".  Its sufficiently neutral to allow both
designs that have "chapter numbers" and those that have no "chapter numbers".
[Within each contribution, it does seem more like a "journal article" than
like a "\chapter" so, if \volumeeditor was adopted to mean "volume editor",
one could probably use \author rather than \contributor without (too much)
confusion.]  Hence, if necessary, \volumetableofcontents,
\contribtableofcontents, etc. perhaps?
If an approach such as that outlined above was adopted, only a small

proportion of the people involved (the volume-editor, the publisher's staff,
the style-file hacker) would need to be aware of any of the above.
The individual contributor to a multi-author work can regard themselves
as producing a paper (e.g. in 2.09's \documentstyle{article});  indeed
the volume-editor might send them some "instructions for contributors" that
tell them to do just that.  So specialised commands/environments (such
as "contribution") needn't necessarily be mentioned in the documentation
aimed at ordinary end-users/contributors.  Contributors could pick up
\section etc. from the body of the LaTeX 3.0 manual; details (of specialised
environments like "contribution") needed by volume-editors could be given
in an appendix of the LaTeX 3.0 manual or just as comments in the style files.

—

From:     David_Rhead@UK.AC.NOTTINGHAM.CCC.VME
Date:     Mon, 15 Apr 91 10:28:13 BST
Michael commented that:
    " ... the way the \appendix command works has always seemed slightly odd
    to me. It would seem more natural to make \appendix a direct substitute
    for \section (in an article documentstyle) or \chapter (in a book
    documentstyle). Maybe in LaTeX 3.0, with its attribute handling, we'll be
    able to say something like \section[variant=appendix]{...}"
I presume that doing the above literally would conflict with the requirement
that LaTeX 3.0 should be able to process LaTeX 2.09 input files (which might
use \appendix as defined in the 2.09 manual).  So \appendix would have
to be left with its 2.09 meaning, but something better could be phased in.
Perhaps one could have
..
\begin{backmatter}
   \begin{appendices}
   ...
   \end{appendices}
   ...
\end{backmatter}
I think that, if a document has just one appendix, then it is silly to call
it "Appendix A".  It can just be called "Appendix" (since there is no
"Appendix B").  If LaTeX can't work out for itself that there is only
one appendix (e.g. from the .aux file), it might be appropriate to allow
a variation of "appendices" that tells LaTeX that there is only one appendix,
e.g. \begin{anappendix} ... \end{anappendix}.
There is the problem of how to introduce the individual appendices:  by using
\chapter (in book, report, etc.) or \section (in article etc.), or by using
something else?  I don't have any specific suggestions here (unless anyone
likes \toplevel), but I think that the question needs considering in

conjunction with questions about how to introduce other "top level" units
that appear in the front-matter and back-matter (such as Preface,
Acknowledgements, Glossary).

—

From:     Michael Downes <MJD@com.ams.math>
Date:     Mon 15 Apr 91 12:55:35-EST
To:       David_Rhead ...
> I think that, if a document has just one appendix, then it is silly to call
> it "Appendix A".  It can just be called "Appendix" (since there is no
> "Appendix B").
I was thinking about this very thing over the weekend. I surmised
that in LaTeX 2.09 the user should be instructed to do something like
       \appendix
       \renewcommand{thechapter}{}
       \chapter{Text of the title}
(in a book). But I haven't tried this yet to see what might happen with
table of contents entries and running heads.
...

—

From:     N.POPPELIER@NL.ELSEVIER
Date:     Mon, 22 Apr 91 08:34:57 +0000
The TEI draft report is certainly easier to read than the AAP documentation.
Furthermore, the AAP dtd is more or less a dtd for the exact sciences, you
know: with lots of math and tables. TEI aims at linguistics, humanities, etc.
It is useful to have a look at both, since two of the flaws of LaTeX's
document styles are, in my opinion,
1. not enough structure
2. aimed at exact sciences (with exception of chemistry),
   hardly usable for humanities

—

From:     N.POPPELIER@NL.ELSEVIER
Date:     Mon, 22 Apr 91 08:48:54 +0000
I am strongly in favor of the nomenclature of the Electronic Manuscript

Standard (ANSI Z39.59-1988), otherwise known as the AAP dtd.
book|article|journal == (fm, bdy, bm)
fm: front matter
bdy: body, which in turn is defined as 1 or more chapters|sections|articles
bm: back matter

—

From:     CHAA006@UK.AC.RHBNC.VAX
Date:     Mon, 22 Apr 91 13:57:27 BST
Bearing in mind Franks' frequent injunctions to make a contribution
rather than allow silence to imply tactit acceptance, I would like
to add my vote to support (front matter, body matter, back matter).
Strangely I would prefer end matter to back matter for reasons which
are not clear even to me, but realise that the term is ambiguous and
could encompass front matter as well as back matter.
     ** Phil.

—

From:     MITTELBACH FRANK <PZF5HZ@EARN.RUIPC1E>
Date:     Tue, 23 Apr 91 15:34:29 CET
I just finished an issue of some internal newletter printed at the EDS
which has the problem of being a typical multi-author work. There is
the overall stuff like table of contents as well as individual articles
from different authors. There is also some `backmatter' material
like a glossary which is compiled by the editor.
The individual articles are of (at the moment) two different types
namely `contributions' and `notices' in our internal jargon.
Guided by production experiences the style for this newsletter
was set up in a way that individual articles are prepared in
the standard article style syntax.
Reasons:
   1) The individual author is able to typeset his/her article
      with any style (house or otherwise) that implements `article
      makeup'. He is not forced to know about special names and
      functions but can view his article in the way the document is
      natural to him, namely as an individual article which is so far
      not necessarily part of some larger structure.

2) The only thing we request is to separate the article from
the driver file to keep unnecessary editorial work to a minimum.
That is, we suggest to use
 \documentstyle{article}
 \begin{document}
   \input{myarticle}
 \end{document}
The file myarticle then contains the real article with
definitions and all, and can therefore be pasted directly into
other structures.
3) The style itself contains environments like `notice' and
`contribution' that redefines \author \maketitle and the like
to produce a desired layout. Some of these commands are ignored
depending on the kind of article. This means that the editor is
free to turn any article into a contribution or a notice or ...
without changing the source.
This concept was considered important as it speeds up processing
time. Changes to articles are done only for the sake of
editorial corrections (typos, grammar, contents) but are not
necessary when simply fitting the article into the larger
context of the newsletter.
4) On the newletter surface, commands like \tableofcontents
are available but produce information assembled from \maketitle
etc of individual contributions and notices.
5) We think of adding a section `letters to the editor' which will
certainly be implemented as some letter environment that accepts
standard letter GML but turns this into suitable output for the
newletter.

Problems:
As mentioned by several people on the list, the front matter
declarations in LaTeX are clearly not sufficiant. In fact they
are nearly non existant. This means that we currently have to
help ourselfs by abusing fields like \date which then in turn
produce a problem in such a concept.
But this only means that we have to define a proper front matter
concept that will allow to specify necessary information in a
natural manner.

  I don't consider the question of compatibility here a difficult one
  since the current LaTeX provides nearly nothing which can be easily
  provided in a `compatibility style' or changed in an older source.
  The use of identical names(e.g. \tableofcontents) for slightly
  different things might be considered confusing but this is something
  I think is in practice perhaps even more natural then inventing new
  names (like \volumetoc \issuetoc ...) that are actually all commands
  for the same concept only that the special outcome is determined by
  the position in the document.

—

From:     MITTELBACH FRANK <PZF5HZ@EARN.RUIPC1E>
Date:     Tue, 23 Apr 91 15:35:34 CET
David said some time ago:
>
> I presume that doing the above literally would conflict with the requirement
> that LaTeX 3.0 should be able to process LaTeX 2.09 input files (which might
> use \appendix as defined in the 2.09 manual).  So \appendix would have
> to be left with its 2.09 meaning, but something better could be phased in.
Let me say a few words to the compatibility problem.
While compatibility is certainly an important issue, we
should not make it the main issue. Seeking better
solutions to some problems might result in some
incompatibility in the end but once we found a solution
we will probably also find a way to successfully parse
older documents.
I think that it is wrong to start with the idea ``LaTeX
2.09 has an \appendix therefore LaTeX 3 needs to have
the same concept.'' To start with the goal that we
``necessarily can process LaTeX 2.09 *within* 3.0
documents without any changes to the old sources''
seems to be too restrictive (to me).  We should try to
achieve this if it is feasible but we should not defend
bad concepts for the sake of full compatibility.
In other words: when we found some concept feasible we
can then ask how to achieve compatibility but not the
other way around.

—

From:     Don Hosek <DHOSEK@EDU.CLAREMONT.HMCVAX>
Date:     Tue, 30 Apr 91 23:34:00 PDT
Some days ago, it had been suggested that the current syntax of
\appendix
\chapter {...}
..
\chapter {...}
..
be replaced with something more like
\appendix{...}
..
\appendix{...}
..
I would like to suggest that this is a mistake. When I teach
LaTeX classes, my presentation of the \appendix command is that
it switches LaTeX into an "appendix mode" where formatting rules
may change. In the standard report style, the numbering of
chapters is changed from arabic numerals to uppercase latin
letters. In other styles more dramatic changes may be called for:
at BRL their article style, in addition to changing section
numbering also changes equation numbering from 1,2,3 through the
document to A1,A2,A3...; B1,B2,B3... in each appendix. The
working design for my class notes switches to nine point type and
two columns for the appendices.
I find that presenting the material in this fashion (and making
analogies with it for the \ps command of the letter style),
allows users to see pass the seeming illogicality of the scheme
and have it make sense again. What's more, it seems to me that
this scheme is somewhat _more_logical in terms of document
structures and also (slightly) easier to implement.

Appendix B

Support for accurate page-selection from dvi files

B.1 Objectives

Although I assume that it is open to anyone writing style-files to do whatever they like with the 10 \count registers [16, p. 119], I think that it would be worth trying to set a good precedent in the L^{A}TEX 3.0 “standard styles”. I therefore suggest that there should be a scheme for using the \count registers in such a way that pages can always be specified unambiguously in those report-like and book-like documents that have:

• roman-numbered “front matter” followed by arabic-numbered body text, and/or

• pages numbered “within chapters”.

In practice, I have the impression that documents which have pages numbered “within chapters” are in a minority. It may therefore be worth having a scheme that can be used selectively, i.e, such that, if a style-file writer knows that it will be important to distinguish between certain units of a document but not between others, the \counts for the former can have the non-zero values given by the scheme, while those for the latter can be zero. This would make for consistency (i.e., if a \count is nonzero, then it would be used in the same way in one standard style-file as in another) without unnecessary clutter (e.g., if all \count registers but \count0 can be set zero without ambiguity, the user would get monitoring of

the form [1] [2] [3] [4] ... rather than of the form [1.1.1] [2.1.1]
[3.1.1] [4.1.1] ... ).

B.2 A scheme

The following scheme would probably achieve the objectives suggested above.

Major division Minor divisions \count0       \count1           \count2
Front matter  —  page-number -1                0
Main text  chapters  page-number chapter-number  0

                                          1, 2, ...

Back matter:  appendices  page-number appendix number 1
appendices  1, 2, ...
Back matter:  glossary,  page-number other-unit  2
other units  bibliography,  number
index, etc.  1, 2, ...

Examples

1. A pamphlet that is numbered in arabic throughout would only use

\count0.

2. A traditional book could have \count1 set to -1 in the front matter and set to 0 thereafter.

3. A manual that is to be numbered in the style of DEC’s VAX/VMS manuals would use the full scheme, i.e.

• \count2 set to 1 within the appendices and set to 2 after the appendices

• \count1 set to -1 in the front matter, set to “chapter number” within the body text, set to “appendix number” or “other-unit number” in the “back matter”.

• \count0 reset at the start of each chapter and each unit of “back matter”.

4. It would be open to someone producing a style-file for a “traditional book” to increment \count1 and \count2 in the way suggested for “a manual” above, even though \count0 would not be reset at the start of each chapter. Although (after the “front matter”) neither \count1 nor \count2 would affect the printed page-numbers, their settings would, for example, enable someone • using DVItoVDU [20] to give a command that means “go to the

start of chapter 2”

• using dviselect [21] to specify “pick out all the pages in chapter 2”.

Thus there would be consistency across most common classes of document regarding the use of non-zero settings for the \count registers, but the number of registers used in a particular style-file could, if desired, be limited to those needed to avoid ambiguous page-specifications at the dvi-file processing stage.

Notes

1. Although the \count registers give 10 possible “dimensions” of pageidentification, Knuth defines yet another “dimension” with the “romannumbered pages have a negative \count0” convention in the “plain” macros [16, p. 252,362]. I guess that the “plain” macros are intended for typesetting a certain class of document (draft papers?), and that the convention is adequate for this class. However, L^{A}TEX is intended to be used for a wide variety of classes of documents, so I think it is better to “start from first principles” and find a scheme that can cope well with the classes of documents for which L^{A}TEX is likely to be used. The scheme suggested in the above table does not aim for compatibility with “plain” since:

• in terms of “logical structure” there doesn’t seem any reason to treat \count0 differently in the “front matter” rather than (for example) in “chapter 2” or in the “back matter”.

• if the end-user learns that “\count1 and \count2 specify the division you are in”, this will be the only concept they need. They can select the 4th page in the “front matter” and page 4 of chapter 5 in a consistent way, e.g., if dvitops [22] is used,

                   dvitops -f 4.-1 root

and

                   dvitops -f 4.5 root

respectively. If L^{A}TEX 3.0 had the additional concept that “romannumbered pages have \count0 negative”, this would be an unnecessary, redundant, concept for the end-user to learn.

2. In terms of “logical structure”, there is a case for treating the “front matter” in precisely the same way as the “back matter”. This would imply the modifications shown below:

Major division Minor divisions  \count0       \count1   \count2
Front matter table-of-contents, page-number 1 -1
list-of-figures,  2
acknowledgements,  3
preface, etc.  4, ...

…

My impression is that, in practice, the scheme shown on page 64 would be adequate since there is generally less “front matter” than “back matter”. End-users would probably prefer to issue a command such as

      dvitops -f 4.-1 root

(as implied by the scheme shown on page 64) to select the 4th page of their “front matter” rather than

      dvitops -f 4.*.-1 root

(as implied by the modified scheme shown on page 66).

3. If desired, the scheme shown on page 64 could be modified to remove the need for \count2 to be non-zero after the appendices. For example, the convention might be:

Major division Minor divisions \count0       \count1   \count2

…

Back matter: glossary page-number -2 0
other units  bibliography

index, etc.

Although this modified scheme would not eliminate ambiguity when people are selecting pages from documents that are numbered like DEC’s VAX/VMS manuals (in which each “other unit” of “back matter” has page-numbers starting from 1), but it might be argued that such manuals are fairly rare.

B.3 Some e-mail comments

From:     bbeeton <BNB@COM.AMS.MATH>
Date:     Thu, 27 Jun 91 11:13:09 CET
re making page numbers distinct, at the american math society, we
faced this problem long ago (though not in latex).  the circumstance
was multi-column pages where each column had to be output separately
on account of memory limitations (this was tex80), and "pasted up"
by the output driver.  the convention we devised (and still use with
some publications) is
  \count0 = \pageno (for compatibility with plain)
  \count1 = column within page (for consistency, if 1-column pages
            were mixed with 2-column pages -- an everyday occurrence --
            two "logical" pages were put in the dvi file for each
            "physical" page)
  \count2 = sequential counter, 1, ..., col x number of pages
although it doesn't necessarily bear any clear relation to the printed
page number, the sequential counter does guarantee a unique value that
can be homed in on by a device driver (although not all device drivers
may give you access to all ten counters, a difficulty that should be
decreasing), and is very nicely reported by tex in the transcript file.
if used in conjunction with some more descriptive scheme, it gives a
useful value that can help one identify pages properly even to a driver
that gives no access at all to the tex counters but always considers
the first page of a dvi file to be page "1" (another driver lapse that
should mercifully soon be a thing of the past).

—

From:     eijkhout@EDU.UIUC.CSRD
Date:     Thu, 27 Jun 91 08:45:10 -0500
David Rhead's problem is a legitimate one. Here's how I have solved
it (for the moment) in my own format:
- use count0 for counting succesful shipouts by the output routine;
- use count1 for keeping track of the number at the top/bottom of the page.
That way the user can let the page number jump any way s/he likes, and
still no dvice driver will ever complain about pages not found,

which could still happen if count0 is merely the page number.
I guess this could be augmented by a facility to print both counters
at the page bottom, but I found it easy enough to read the correspondence
from the log file (the joys of X windows: there is always some corner left
on the screen to display some piece of information).
Victor Eijkhout

Appendix C

Some e-mail comments about treatment of wide objects

From:     "Nelson H.F. Beebe" <beebe@EDU.UTAH.MATH>
Date:     Tue, 18 Dec 90 17:17:39 MST
In the current LaTeX, input of the form
\begin{center}
...something...
\end{center}
does not center the ...something... if its box is wider than
textwidth; instead, it is flush against the left margin, and
extends into the right margin.
I feel this is counterintuitive, and the situation happens
often enough with wide tables, that I sought, and found, a
solution:
% This environment allows wide centered objects to spill into the
% margins, evenly on both sides
\newenvironment{widecenter}{
    \bgroup
    \makeatletter
    \@flushglue = 0pt plus 1fil minus \textwidth
    \makeatother
    \begin{center}
}{
    \end{center}
    \egroup
}

This solves my own problems, but the question for the LaTeX
redesign is:
        Should centering into the margins be the default
        behavior of the center environment?
Obviously, the default definition of the center environment could
be easily modified to work this way.
I personally would vote for such a change, unless someone exhibits
good reasons to the contrary.

—

From:     bbeeton <BNB@COM.AMS.MATH>
Date:     Thu, 20 Dec 90 16:32:56 CET
nelson asks whether over-large centered items should be allowed
to exceed the margins automatically.
while i agree that centering of items that are slightly wider
than the text area is usually preferable to having it shifted
to the right, in cases such as centered titles an even better
solution is to divide the centered text into more shorter lines.
it's my feeling that many people tend to ignore overfull boxes in
any event, and full automatic centering might tend to camouflage
situations where reasonable action would yield a better result.
so i'd be in favor of keeping a widecenter environment, for
intentional use in situations where there isn't any choice.

—

From:     MITTELBACH FRANK <PZF5HZ@EARN.RUIPC1E>
Date:     Thu, 20 Dec 90 17:46:11 CET
>
> while i agree that centering of items that are slightly wider
> than the text area is usually preferable to having it shifted
> to the right, in cases such as centered titles an even better
> solution is to divide the centered text into more shorter lines.
>
> it's my feeling that many people tend to ignore overfull boxes in
> any event, and full automatic centering might tend to camouflage
> situations where reasonable action would yield a better result.
> so i'd be in favor of keeping a widecenter environment, for
> intentional use in situations where there isn't any choice.

What Barbara is implicitly pointing out is that there are two
or perhaps three distinct concepts connected with centering of
items.
  1) Centering of text and objects similar to text in a
     `reasonable' way. If things do not fit into a
     given frame (\hsize) then the user gets a waring
     and an overfull box.  This is more or less
     realized in the currrent LaTeX environment.
  2) Centering of items on the current page regardless of their
     size, or at least within some given extra size. This is
     useful for tables figures etc. and only possible in the
     current LaTeX by using \makebox to hide the actual size.
  3) Centering of text within a given frame so that lines are
     neither too long nor too short. This can be done automatically
     only making direct use of \leftskip and \rightskip.
I suppose the correct way to handle this is to either provide
different environment or (which I favor) to add attributes to the
center environment to address the special cases. I think that the
default setting could be left as it is.
A similar problem arises with flushleft and flushright environments.
In case of flushleft we miss a raggedright environment for text
typesetting which is able to hyphenate in a given range.

—

Subject:  Re: Wide things
From:     Don Hosek <DHOSEK@EDU.CLAREMONT.HMCVAX>
>My first reservation is that, in publishing practice, it seems normal
>[1, p. 336; 2, p. 158] to set tables in a smaller typeface than the body of
>the text.  [1] says "tables may be set in small type or footnote size, or the
>size may vary from one table to another ... ".  [2] seems to imply that it
>is normal to set tables in an 8-point typeface, presumably when the
>body of a document is in 10-point (and suggests going down to 7- or 6-point
>if a table won't fit in 8-point).  In fact [1] and [2] recognise that
>"getting tables to fit" is often difficult, and devote several paragraphs
>to tactics for dealing with the difficulties.  So I think that:
>*  the user should still get a warning if a centred thing is allowed to
>   spill into the margin.  The user should then e.g. go and read [1] and [2]
>   to decide whether to accept the "spillage" or to e.g. see whether
>   column-headings could be made more concise and/or decide whether to change
>   font-size

>*  there is a case for making the font-size in the "successor to tabular"
>   a feature of the design.  Typically if a design has main text in 10-point,
>   the design might give 8-point as the default in the "successor to tabular".
>   Then (a) LaTeX would conform to what appears to me to be publishing
>   practice (b) the proportion of tables whose width causes problems would
>   be reduced (and less users would bother people like me about them).
Below is an excerpt from my LaTeX class notes which explains how
I teach users to deal with the situation here and now. Although the
text deals simply with centering and the like, the approach can be
applied to other items of table formatting as well (many of my LaTeX
styles provide a somewhat fancier aroundtbl environment built-in).
The tabular environment is normally treated as just a big letter.
For example, it's possible to get LaTeX to typeset something that
            4  1    ii
looks like 123 345 fgh in the middle of a paragraph. This is
            7  12    q
seldom actually desired, but useful to know. Where this becomes
particularly noticable is that a tabular environment in its own
paragraph is typeset with paragraph indentation, viz:
    4  1    ii
   123 345 fgh
    7  12    q
rather than flush left as you might prefer. The best way to deal
with this problem is to define an aroundtbl environment which
will tell LaTeX how to postion the tabular. The reason for
defining an environment in this fashion is that it makes it
easier to change how the tabulars will be placed: if one simply
told LaTeX to center each tabular and then decided to typeset all
tabulars flushleft, every instance of tabular would need to be
changed. Defining the aroundtbl environment would reduce the work
to simply changing a single definition.
Using the aroundtbl environment, one would simply type something
along the lines of the following for each tabular:
\begin{aroundtbl}
\begin{tabular}{cc}
These columns & are \\
lined up      & and centered
\end{tabular}
\end{aroundtbl}

The following definitions are three possibilities for definining
the aroundtbl environment:
\newenvironment{aroundtbl}{\begin{center}}{\end{center}}
will center each tabular.
\newenvironment{aroundtbl}{\begin{flushright}}{\end{flushright}}
will place each tabular flush against the right margin.
\newenvironment{aroundtbl}{\begin{flushleft}}{\end{flushleft}}
will place each tabular flush against the left margin.
n.b. You should never use the center, flushright, or flushleft
environments outside of the definition of a new command or
environment.

—

From:     Leslie Lamport <lamport@COM.DEC.SRC>
Date:     Wed, 2 Jan 91 20:12:58 CET
Once again, the discussion on LaTeX-L is sailing off into the wild blue
yonder.  The document style should not try to figure out that because
this is the left-hand page of an article for a physics journal edited
by a woman, overlong equations should extend to the right.  This kind
of nonsense will produce 126-megabyte document styles that will exhibit
three new bugs per month for ten years.
Problems such as things that don't fit on a line require user
intervention.  They are solved, as the final step in producing a
document, by inserting "visual-formating" commands.
Instead of all this pie-in-the-sky discussion about how to do
everything automagically, how about giving some thought to how the user
can EASILY specify where this particular equation or table should be
placed on this page of this version of this document.  Or, more
generally, on how visual-formating commands, which may have to be
changed each time a new version is created, should be integrated
with the idealogically correct logical-formating commands.

—

From:     David Rhead ...
Subject:  Tables (which may be wide)

Date:     9 Jan 91 15:25:49
To:       latex-l%dhdurz1@uk.ac.earn-relay
Msg ID:   < 9 Jan 91 15:25:49 GMT A101BA@UK.AC.NOTT.VME>
Would there be a consensus in favour of the following?
1.  Have a standard name for "an environment into which people normally
put non-floated tables".  Names currently on offer include Don
Hosek's "aroundtbl" and my "displaytable".  I'll use the name "N"
below (where N denotes an unknown).
2.  In the manual, mention N (as the environment into which the user
normally puts a table) before mentioning center, etc. (to which the user
might have to resort if user intervention is required for a particularly
awkwardly sized table).
3.  In the standard styles, have a fairly simple definition of N, but
one which conforms to a common publishing convention. E.g., in the
successors to report, article and book, something like
   \newenvironment{N}{\begin{center}\begin{footnotesize}}%
   {\end{footnotesize}\end{center}}
would do (although a definition involving Nelson's widecenter
or Frank's "center with attributes" could be used). Thus the user
would go
  \begin{N}
  \begin{tabular}{...}  %  or tabbing, or their successors
  ...
  \end{tabular}
  \end{N}
I'd justify the footnotesize as a "common publishing convention" by
reference to Butcher's book and the Chicago Manual of Style (see my last
mail) which both talk in terms of tables usually being set in a smaller
typeface than the main text (but people who work for publishers would
be better placed than me to advise).  Such definitions should be
sufficiently simple that they don't generate "3 bugs/month for 10 years".
Yet the existence of N would make it clear "what one should do with a
table" and having a reduction in size as standard (if adopted) would
save a bit of "support" time by allowing more information in before
problems of size arise.
4.  If someone is writing a style-file for a particular application, and has
to implement a particular design, it will be up to them how ambitious
they are in their re-definition of N.
People attenting Don Hosek's classes might simply replace center by
flushright or flushleft.

If a "a woman editor of a physics journal" has to enforce a house-style
that doesn't lend itself to a simple re-definition of N, it would be up
to her whether she goes for:
(a) a complicated re-definition of N, in the hope that, although she can
   expect to spend time finding bugs in her definition, the time she spends
   on coding and bug-fixes may be less than the time she would have spent
   in "visual formatting",
(b) visual formatting, by replacing N by lower-level commands in the
   few cases when the simple definition leaves things sticking into
   a margin (or margins).
If she goes for (a), and is sufficiently confident of the code to put it
"in the public domain", people will look to her for maintenance of the
code, not to the people who produced the simple code in the standard
styles.
Similarly, it would be open for someone writing their own style file
to try making N intelligent enough to "fit" awkwardly sized tables
automatically, but maintenance would be a matter for them.  E.g. if
someone produces a style-file that lays a document out in the style
of the PostScript manuals, it would be up to them whether they try
to make N intelligent enough to use the "heading margin" when "necessary",
and if so whether they risk putting the style-file in the public domain.
Nothing they did would imply any additional maintenance for the
simple definitions used in the "standard styles".

There would then be the following open questions:
*  the name to be used for N
*  the relationship with the "floating environments".
Should it be possible to have a common numbering sequence for the captions
for both fixed and floating tables?  If so, it would presumably be necessary
to define N so that, as far as floats were concerned, it was as if a
\begin{table}[h] had been possible.  Otherwise, numbering could get out of
sequence if a floated-table floated past a fixed-table.
Should a float automatically do an N?  If it's an environment that's
specifically designed to float tables, I'd say "yes" for the standard
styles, although obviously anyone writing anything else could do what
they liked.

—

From:     MITTELBACH FRANK <PZF5HZ@EARN.RUIPC1E>
Date:     Thu, 10 Jan 91 14:11:27 CET
>

> Once again, the discussion on LaTeX-L is sailing off into the wild blue
> yonder.  The document style should not try to figure out that because
> this is the left-hand page of an article for a physics journal edited
> by a woman, overlong equations should extend to the right.  This kind
> of nonsense will produce 126-megabyte document styles that will exhibit
> three new bugs per month for ten years.
Of course, the document style should be able to figure
out that this is a left hand page and place an example
(as in the LaTeX book) into the correct margin. The
same might be true for other objects in special
application. This is not a question of standard styles,
but one of my goals for the LaTeX reimplementation is
to provide the necessary tools for the designer to
produce such styles if desired.
...
In this special case, the concept of a
`displaysomething' environment looks like a good idea
to me, at least much better then telling the user to
use center, etc. for things like this. ...
>
> Problems such as things that don't fit on a line require user
> intervention.  They are solved, as the final step in producing a
> document, by inserting "visual-formating" commands.
>
That there are a lot of cases which can not be solved
automatically is clear (for this reason I like to have
e.g. user/editor invention via attribute). But is not a
bad idea to free a standard document from  visual
formatting by adding standard logical concepts which
are missing so far (do not let you guide solely by
trying to keep the LaTeX book rewrite to a minimum).
This can and should be discussed, that just the reason
for this list.
> placed on this page of this version of this document.  Or, more
> generally, on how visual-formating commands, which may have to be
> changed each time a new version is created, should be integrated
> with the idealogically correct logical-formating commands.
>
This is an interesting side remark, should there be
some way to make visual formatting like \newpage
visible? I remember a case where somebody said
\samepage without any grouping and it took me a long

time to find out why his document was formatted in a
funny way.

Appendix D

Citations and reference-lists

D.1 Some e-mail comments

From:     Sebastian Rahtz <spqr@UK.AC.SOTON.ECS>
Date:     Tue, 7 Aug 90 15:33:13 bst
...  writes:
 > One of the most common mistakes that I see from LaTeX users is typing
 > \maketitle before \begin{document}; since this causes horrible things to
 > happen, perhaps we should make sure that this causes an error.
and why not put \bibliographystyle before \begin{document}, where it
belongs..

—

From      David Rhead ...
Date:     13 Aug 90 17:39:39
Sebastian suggested
     "and why not put \bibliographystyle before \begin{document}, where it
      belongs.."
This might need thinking through, bearing in mind that certain types of
document may have more than one list of references.  For example:
Manuals produced by software houses, e.g. the SPSS-X documentation.
      Such manuals may, in effect, divide citations into 2 categories:
      1.  references to other manuals produced by the software house,
          which might be cited using a "short title" scheme
      2.  references to other literature, which might be cited using
          an author-date scheme.
      E.g. the SPSS-X Introductory Statistics Guide generally uses
      author-date, but gives the full reference to the SPSS-X User Guide

in its preface and thereafter refers to it as "SPSS-X Users Guide".
The software house may list its own publications in the preface to
its manual, putting the list of references to other literature
at the back of the manual.  To support this sort of thing would
require something like
      \documentstyle{manual}
      \begin{document}
      ...
      %  \xxx represents a command that embeds \bibliography-like
      %  information in a preface.  "software-house" represents
      %  a style that software houses seem to like where the
      %  reference is embedded in an explanatory paragraph.
      \xxx[Our other manuals]{software-house}{our-manuals}
      \chapter{...}
      ...
      \bibliography[Other people's stuff]{author-date}{other-literature}
where \xxx and \bibliography are assumed to have
*  an optional argument to specify a title
*  an argument to specify the scheme (like \bibliographystyle does)
*  an argument to specify the bib files.
Books.  It might be sensible to divide a book's references into 2 lists, e.g.
"References" and "Further reading".  E.g. the draft revised British
Standard for theses suggests having "Bibliography" as well as "List
of references".  It might also be sensible to have different styles
for the 2 lists:  perhaps a concise style for the "References",
but a style that prints some extra information about the
"Further reading".  The author might want to specify something like
      \documentstyle{...}
      \begin{document}
      ...
      \bibliography[References]{author-date}{specific-books}
      \bibliography[Further reading]{annotated}{general-books}
      ...
[Such sub-division seems to be countenanced by the gurus, e.g. Chicago
Manual of Style (p. 425) and Butcher's Copy-editing (p. 183,192).]
Unfortunately, such considerations lead to more questions like "How does
one arrange that \cite gives (...) around author-date citations, but not around
short-title citations?"  and "Can a root file have several bbl files and what
would they be called?", and I don't know the answers.
At the moment, I just think that it would be better to refrain from putting
\bibliographystyle before \begin{document} until the implications for
documents that have more than one list of references have been thought through.
Otherwise there might be a change to "\bibliographystyle before
\begin{document}" at LaTeX 3.0 followed by another change (e.g. to
"bibliography style as argument of \bibliography") in some subsequent LaTeX.

—

From:     Sebastian Rahtz <spqr@UK.AC.SOTON.ECS>
Date:     Tue, 14 Aug 90 10:00:33 bst
 >      "and why not put \bibliographystyle before \begin{document}, where it
 >       belongs.."
David Rhead has some sensible cautionary remarks, and I would concur
with his underlying thesis that the bibliography support in
LaTeX/BibTeX has definite lacunae. He identifies the problem
 > This might need thinking through, bearing in mind that certain types of
 > document may have more than one list of references.  For example:
and suggests a formulation along the lines of
 >             \bibliography[References]{author-date}{specific-books}
 >             \bibliography[Further reading]{annotated}{general-books}
this is interesting, but nobody will thank anyone for changing the
basic syntax of a LaTeX command in the near future. I would suggest
retaining \bibliography as it is documented, and implementing David's
suggestions with others as new commands. ... such as separate
bibliographies for chapters - whatever LL or OP say about it being so much
work to do a multi-chapter book that sets of bibliographies being not
much more work, I don't see why we shouldn't get it in one day. People
*do* want it. I am on my fourth conference proceedings in as many
years, and I dont enjoy building the slightly complicated Makefile to
get all the references in the right place up to date.
but I continue to say that the user who puts \bibliographystyle before
\begin{document} is behaving intuitively, and should therefore be
allowed to do it. lets keep `bibliography' for the relatively simple
academic type that LL and OP envisaged, and invent a new term
(\reference?) for what David Rhead is talking about.

—

From:     "Nelson H. F. Beebe" <Beebe@EDU.UTAH.SCIENCE>
Date:     Sat, 18 Aug 90 17:15:10 CET
I just hit what I view as a design flaw in thebibliography
environment.  It was precipitated by the following BibTeX
bibliography entry in the .bbl file:
\bibitem{Steele:floating-point-output}
Guy~L. {Steele Jr.} and Jon~L. White.
\newblock How to print floating-point numbers accurately.
\newblock {\em ACM SIG\-PLAN Notices}, 25(6):112--126, June 1990.
\newblock In electronic mail dated Wed, 27 Jun 90 11:55:36 EDT, Guy
Steele reported that an intrepid pre-SIGPLAN 90 conference

implementation of what is stated in the paper revealed 3 mistakes:
  \begin{enumerate}
    \item
         Table~9 (page 125):\par\noindent
        \begin{tabular}{ll}
          for & {\tt -1:USER!({"}{"});} \\
          substitute & {\tt -1:USER!({"}0{"});}
        \end{tabular} \par\noindent
       and delete the comment.
    \item
          Table~10 (page 125):\par\noindent
          \begin{tabular}{ll}
            for & {\tt fill(-k, {"}0{"})}\\
            substitute & {\tt fill(-k-1, {"}0{"})}
          \end{tabular}
    \item
          Table~5 (page 124):\par\noindent
          insert {\tt k <-- 0} after assertion, and also delete {\tt k
          <-- 0} from Table~6.
  \end{enumerate}
The effect of this is that the next bibliography entry gets
number 4, one more than the last enumerate counter.  The reason
lies in latex.tex:
%  The thebibliography environment is a list environment.  To save the
%  use of an extra counter, it should use  enumiv  as the item counter.
For now, I will switch to an itemize instead of enumerate.
The question for this list is, is the saving of an extra counter
at what is usually almost the end of a document anyway worth this
design gotcha?  I suggest not.

—

From:     MITTELBACH FRANK <PZF5HZ@EARN.DRUEDS2>
Date:     Fri, 5 Oct 90 17:43:46 CET
I would like to foward a mail from Nico which was a reaction
to some discussions at the Cork meeting.
------------------------ forwarded mail -------------
>
> here are some first thoughts on BibTeX and front matter.
>
> BibTeX

> ------
> BibTeX is unfortunately not a database management system --
> unfortunate in the sense that users have to think of a way to
> *really* manage the database, i.e. add, change, remove, sort, select,
> import and export entries. On my Atari I use a simple database
> program -- a real GEM program with mouse and buttons -- to maintain a
> literature database. I've written a conversion program that generates
> a .bib file from this database. This system works and is also used ny
> my ex-colleagues at the university.
>
> In general, working with BibTeX is cumbersome -- every time you add a
> reference to your document you have to run LaTeX, BibTeX, and LaTeX
> twice again -- and a lot of LaTeX users think it only pays off when
> your document/bibliography is above a certain length.
>
> What I would like to see in LaTeX is coding of the logical structure
> of every \bibitem. In version 2.09, BibTeX takes care of the logical
> structure and outputs formatted text. In other words: a part of the
> document, the one included by the \bibliography command, does not
> contain tags for the logical structure. Furthermore, if a user
> decides NOT to use BibTeX, he/she has to do the formatting completely
> by hand.
>
> Front matter
> ------------
> Currently, an article starts with \title, \author and \date
> instructions. In the Elsevier styles I have added \address,
> \received, \revised and \accepted commands, and also a keyword
> environment, similar to the abstract environment. By doing so, we can
> automatically convert a LaTeX-coded document to an SGML-coded
> document.
> I think it is a bad idea to put the front matter information in a
> .bib record for the above reasons and also for the following reason:
> some parts of the front matter do not have to be re-used again in
> bibliographical entries, and therefore do not belong in the .bib
> database.
> Examples: list pf previous books in a series, LCC data, dedication,
> motto (books), keywords, date of receipt, revision, acceptance
> (articles).
>
...
The last comment about the bad idea probably needs some
explanation: During the Cork conference we discussed the possibility
to use an extended version of bibtex, which is able to produce
several bibliogrphies at once, to format a title page (on request).

The idea was that for larger documents one could set up a bibliography
entry for the document itself with all relevant information in it.
In the document itself the title would then be formated by a command,
say \titlefrombib{<key>} which would make request bibtex to produce
a file containing all necessary information for producing the title.
This means that every journal that accepts LaTeX input would supply
a BibTeX style file which generates from a single bib entry such
a file. The advantages of this method would be that unused information
would produce no problem (the bst file would simply ignore them).
Another advantage would be that authors (for eample for TUGboat) would
already sending a bib entry along with his document.
Of course, a standard way for simple document that generates titles
without using BibTeX should be availabe too. And the standard document
styles should also have their bst files for producing the title.
This scheme if, of course only sensible (if at all) if \BibTeX is
able to produce several bbl files in one pass.
I still think that such a feature has some promising possibilites
and would like to hear other opinions about it.

—

From:     David Rhead
Date:     9 Oct 90 10:39:45
Frank forwarded some mail from Nico about BibTeX etc.

BIBLIOGRAPHIC DATABASE?
I agree with Nico's comment about BibTeX not being a database management
system.  I have the impression that some research workers really want a
system that will not only contain author, date, etc. for a paper/book,
but also a copy of the abstract and perhaps some reminders to themselves.
Thus, they want something that will both
(1) help (by supplying bibliographic details) when they are in the process
    of writing a paper
and
(2) help them to search through all the papers they've ever read
    until they come to something that is about a particular subject.
    E.g. "give me all the papers about X", "what was that paper about Y?".
It's unfortunate that Lamport/Patashnik used the term "bibliographic
database" (although I can't think of anything better offhand).  It leads
to people thinking that BibTeX will do the things that they associate
with "database systems" these days.

However, I think that to turn BibTeX into something that did more
than (1) would be too ambitious.  We're going to have enough problems
finding someone to make BibTeX do (1) better.
I think it would be better to treat (2) as a separate project, and to ask
"is (2) best done by adding database management features to BibTeX,
or would it be better done by adding BibTeX-like features to a
database management system?"  I don't know what the answer is.
For the time being, I think users will have to be left in an unsatisfactory
situation.  If the "raw data" is kept in a "real database", the .bib file
is just yet another intermediate file.  I seem to remember that
Sebastian Rahtz has set INGRES up so that INGRES can hold bibliographic
information and write it out in the form of a .bib file.  It sounds as
though Nico is doing something similar with a different database
system.
It seems sensible to use a "real database system" for what "real database
systems" are good at.  Although it is clearly unsatisfactory to have a .bib
file that merely "shadows" a "real database" I can't think of anything better
that could be done quickly.  [Perhaps a database expert might have some bright
ideas.  Could such an expert write software that took an .aux file, generated
instructions in a "query language" to select the \cite-ed references, and
then produced a .bbl file (or equivalent) without there ever being a .bib file?
I don't know:  I'm not a database expert.  Even if they could, it
would probably still be useful to have a standalone program like BibTeX
that did task (1) in a database-system-independent way.]
A salesman has sent me a leaflet about a piece of software called
EndNote which apparently seems to aim to do both (1) and (2) for Mac and PC
word-processor users.  (It's not public domain, but then neither is INGRES.)
I see that it can export information for troff's "refer".  He's coming to
see me at the beginning of November, so I'll ask if EndNote might be tailored
to read .aux files and produce .bbl files.  I expect the answer will be
"no" or "only after a lot of work", but it does no harm to ask!
Perhaps one could have a public-domain BibTeX for task (1) with tailored
proprietary software for people who want (1)+(2).
Conclusion:  I can't think of an easy way of improving the "3 times
through LaTeX and once through BibTeX" business at the moment.  I doubt
whether its worth making BibTeX into a proper database system, although
it might be worth making "a proper database system" do what BibTeX does.

CODING THE LOGICAL STRUCTURE OF EACH \BIBITEM?

I can see the attraction of this.  The list of references could be
held in the root file or \input or \include-ed from a .tex file
like everything else.  There wouldn't need to be any special treatment
of .bbl files.
However, in terms of project-management, it seems very convenient to regard
determination of the logical structure for bibliographic references
as a separate task which can be delegated to whoever volunteers
(if we can find them) to do another iteration on BibTeX.
It would be up to this person to specify the logical elements
for a reference, e.g. to agonise about whether ADDRESS is
the fundamental concept of whether it should be PLACEOFPUBLICATION.
All that needs agreeing between the LaTeX 3.0 people and the
BibTeX worker (if we find one) is the form of interface, i.e.
what BibTeX passes back to (a) go where the \cite was and (b) go
in the list of references.  There might need to be separate
interfaces for "reference by number", author-date and "short title"
[I'll have a go at suggesting what these might be in a future message]
but beyond that, the LaTeX 3.0 people need not be concerned
about the distinction between e.g. different author-date styles.
Thus, the BibTeX worker's considerations of matters such as
ADDRESS versus PLACEOFPUBLICATION can proceed in parallel with
the LaTeX 3.0 people's work on other matters.  It wouldn't matter
much whether both projects were complete at the same time.
If the LaTeX 3.0 people attempted to code the logical structure of
each \bibitem, this would mean that LaTeX 3.0 could not be finished
until the LaTeX 3.0 people had satisfied themselves about "What are
the fundamental types of publication (\bibitem{knuth-84}{book} or
\bibitem{knuth-84}{monograph}?), and what are the fundamental
items of bibliographic information about them?  Did Lamport/Patashnik
get it right?  E.g. ADDRESS versus PLACEOFPUBLICATION."  The LaTeX 3.0
people would presumably have to provide TeX code that sorted bibliographic
details out into the order required for a particular style e.g. "reference
by number in ACM style" (and perhaps provide a few style-options to
show how the details could be changed for a different convention,
e.g. author-date in APA style).  I think that the LaTeX 3.0 work
is ambitious enough without taking this analysis on too.  Attempting
to get the subdivision of \bibitem right could hold the rest of the
project up.
Continuing to delegate the work on the logical structure of each \bibitem
to BibTeX might not be as elegant from the user's point-of-view as
getting LaTeX to do all the work (using structure information from
subdivisions of \bibitem and bibliography-style information from
\documentstyle) but I think the result would be available sooner and that

the user might prefer to have something better than LaTeX 2.09 soon, rather
than to have perfection not so soon.
The two approaches (1) "put all the BibTeX work into LaTeX" (to take
account of the logical structure of each \bibitem), and (2) "make
BibTeX into a proper database management system" seem to be pulling
in different directions.  I don't think one can do both (otherwise
you'd end up with LaTeX as a bibliographic database management system),
although one could do neither.
For LaTeX 3.0, I'd be inclined to leave the contents of each \bibitem
(or the successor to \bibitem) as a "black box", to be filled in by the
user or by BibTeX.  If someone does the analysis for BibTeX 2.0 (say),
the question could be considered again if there is ever a LaTeX 4.0 (!)
[This all assumes that the SGML-ers have not analysed the structure
of a list of references and hence that someone has to do the analysis.
If the SGML-ers have done the anlysis (for a DTD, perhaps), could they
publish it?]
Conclusion:  I'd specify an interface between LaTeX 3.0 and BibTeX
that would support the "reference by number", author-date and (if
possible) "short title" schemes, but delegate the task of supplying
\bibitems (or whatever) to that specification to whoever updates BibTeX
and its .bst files.

FRONT MATTER INFORMATION IN A .BIB RECORD?
The gurus of "how to do a list of references" seem to agree that
bibliographic details should be as they appear on the title page of
the article, book, etc.  But there are many caveats:
*  Several books by the same author in one bibliography should
   follow the same style (Chicago Manual of Style, p. 441).
*  There are potential problems with names like Tchaikovsky,
   which may appear in different forms (Chaikovsky) on different
   title pages, even though the works are all by the same person.
   [British Standard BS 1629, p. 5]
*  The part of the name not on the title page may be enclosed
   in square brackets (Chicago, p. 441).
*  If the name on the title page is a pseudonym, the author's
   real name may be given in the bibliography in square brackets.
   (Chicago, p. 442).
*  Capitalization, punctuation, etc. of a title may be differ in a
   bibliography from the conventions on the title page (Chicago, p. 447).
   Similarly, compulsory line-breaks may be wanted on the title-page
   but not in the list of references (\\, Lamport's book, p. 164).

   For another example, consider "LaTeX: A Document Preparation System":
   that's not how it appears on the title page.
*  It may be necessary to use discretion about whether to regard
   a subtitle as part of a title or to abbreviate a long title
   (ISO standard 690, p. 5).
*  A bibliography may give "place of publication" in a form that is
   different to that on the title page, using discretion about:
   - whether to list all places where the publisher has offices
     or just one place
   - whether to give further information (if the place of publication
     is not widely known or could be ambiguous).
   (Chicago, p. 456).
*  A bibliography might give a publisher's name in a form that differs
   slightly from that shown on the title page (Chicago, p. 458).
I think that having a BibTeX that can produce several bibliographies at
once would be "a good thing".  For example:
*  for conference proceedings where each contribution may have its own
   list of references
*  for books that may have e.g. "References" and "Further reading"
*  things like the SPSS manual (and other things produced by software
   houses), which seem to give the software-house's related publications
   in a preface, but put "academic references" at the end.
However, I don't necessarily think that the same mechanism should
be used to "derive a publication's title-page from its .bib entry".
Traditionally, bibliography entries have been derived from title-pages
(with some human discretion), rather than the other way round, so its
probably safer to have software that imitates the tradition.  One might
think of having things like
\begin{titlepage}
   \author[bibliography-version]{titlepage-version}
   \title[bibliography-version]{titlepage-version}
   \place[bibliography-version]{titlepage-version}
   \publisher[bibliography-version]{titlepage-version}
\end{titlepage}
\begin{copyrightpage}
   \copyrightholder{...}
   \isbn{...}
\end{copyrightpage}
in the .tex file (where the optional arguments allow humans to exercise
the discretion recommended by the gurus) and having LaTeX produce
perhaps a .bibitem file that the user can append to a suitable .bib file
(maybe after exercising a bit more discretion).  LaTeX could put
information that might conceivably be used by a bibliographer in the
.bibitem file, but refrain from putting information that no bibliographer
would ever want to that file.  TUGboat would get a .bib entry

with each article (but it would be derived automatically from the
article-heading, rather than used to automatically produce the
article-heading).
[Conversely, if the Cork suggestion was adopted, and title-page
information was produced by some future BibTeX from a .bib file, there
would have to be some mechanism to allow for minor variations, e.g.
      TITLE = "LaTeX: A Document Preparation System",
      TITLEPAGETITLE = "LaTeX\\A Document Preparation System"]
The problems that I mentioned in the context of "logical structure
of \bibitem" arise here too.  To write .sty files for (say) book, report,
conference-proceedings and article, you only need to be clear about those
categories (as well as being a TeX wizard and having a lot of time,
perseverance and patience).  To write .bst files with entry-types of book,
report, conference-proceedings, article, you need to be clear whether
they are distinct entry-types or not.  [ISO 690 could be interpreted
as lumping books, reports and conference-proceedings all together
as "monographs".]   You also need to be clear about the fields:
e.g. PLACE or ADDRESS.  So the LaTeX 3.0 code for title-pages would
get held up (and hence LaTeX 3.0 as a whole would get held up)
while someone analysed "the structure of a \bibitem".
On the other hand, if LaTeX 3.0 wrote out a .bibitem file that wasn't
quite what some new BibTeX expected, it wouldn't matter very much,
and could be corrected once it was clear what was required.
Conclusion:  I'd like BibTeX to support multiple lists-of-references,
but think that "LaTeX producing .bib info from titlepage info"
might be better than "BibTeX producing titlepage info from .bib info".

—

D.2 My e-mailed “paper”

D.2.1 Introduction

Since there are now plans for a new version of L^{A}TEX [1, 2], this may be a good time to consider how a future version should deal with citations and reference-lists.

My viewpoint is that of an advisor to authors who use L^{A}TEX 2.09. As such, I’m often in the position of having to decide whether L^{A}TEX acts inappropriately or whether an author is asking for something inappropriate. Generally, I have the impression that L^{A}TEX 2.09 sometimes makes it unnecessarily difficult for people to comply with the conventions that are standard in “academic publishing”.

In making the suggestions that follow, I do realise that it’s easier for me to suggest that the volunteers developing L^{A}TEX 3.0 might provide some new facilities than it is for a volunteer to find time to do the work!

I assume that the reader is familiar with the relevant sections of [1].

D.2.2 Conventions to be supported

Citation schemes

In mainstream publishing [23, 24, 4, 3, 25, 26, 27, 28, 29, 30], there seem to be three basic schemes for citations and the corresponding reference-lists:

reference by number In this scheme, citations are normally numbered in order of first citation. In particular, “order of first citation” is used by over 300 biomedical journals [26, 27], and is specified in the ISO standard [28]. (Thus, in BibTEX terms, the usual numbering sequence is that given by unsrt.) The number is used as a “label” in the reference-list.

author-date There are two main forms of citation, depending on whether or not the cited author’s name occurs naturally in a sentence. In the first case, the citation is of the form “ … (1972) … ”, whereas in the second case it is of the form “ … (Crane, 1972) … ”. There are no “labels” in the reference-list, which is arranged in alphabetical order of authors’ surnames (with supplementary rules for “tie-breaking”).

short form The “short form” scheme is often used when citations occur in footnotes. Typically, the first citation (or the first citation in the current chapter) gives a fairly full reference but subsequent citations use a short form. The “short form” may be introduced within the first citation, or given in a table of abbreviations. There are no “labels” in the reference-list, which may be subdivided by “type of cited document”. The scheme is common in the humanities, but also seems to be used by some software-houses when referring to their own publications (see, for example, [31]).

I think that, in order to make it straightforward to achieve, with L^{A}TEX, the effects that people routinely achieve with traditional publishing procedures, it is desirable that L^{A}TEX should provide explicit support for all three citation schemes.

Notice that:

• It is not, in general, possible to convert a document from one scheme to another (e.g., from “reference by number” to author-date) automatically. Some re-writing is required.

• The number of items of information that need to be available differ between the schemes. For a “reference by number” citation, it is only necessary to keep track of one item (the number), whereas for the other schemes it is necessary to keep track of more than one item (i.e., the author and the date, or the “fairly full form” and the “short form”) so that they can be used separately.

• Occasionally, two different schemes may be used in parallel within the same document. (For example, in [31], a software-house seems to use “short form” when citing its own publications, but author-date when citing other publications.)

Thus, it seems to me that the three schemes are best regarded as logically distinct.

However, within a particular scheme, there are variations of punctuation, etc., that can be regarded as matters of “house style”. For example, some journals that have adopted a “reference by number” scheme use bracketed numerals for citations, while others use superscripts. Such variations can be accomodated by differences between style files.

Additional references

In addition to a list of “works cited”, some documents have an additional reference-list that specifies “further reading” or “all sources consulted” (see [4, pp. 182 & 192], [26, p. 40] and [32, p. 22]). There needs to be some provision for typesetting such additional lists.

D.2.3 Deficiencies in L^{A}TEX 2.09

Citations

When used in conjunction with BibTEX and \bibliographystyle{unsrt}, L^{A}TEX 2.09 makes a good job of organising citations (and sorting the corresponding reference-list) according to the “reference by number” scheme as required by many journals.

Generally, however, L^{A}TEX 2.09 does not provide the breadth or depth of facilities needed to support the variety of requirements for mainstream publishing:

• It is unfortunate that the “reference by number” sequence obtained most naturally by the do-it-yourself-er [1, p. 73] from L^{A}TEX 2.09 (“order of appearance within thebibliography”) is unlikely to be what the do-it-yourself-er’s journal-editor requires (which will usually be “order of first citation”). The do-it-yourself-er is given no warning (either in [1] or by the software) that sorting is likely to be needed.

• Although one can use style-options such as apalike and aaai to redefine \cite and thebibliography for an author-date scheme, there are many obstacles placed in the user’s way:

– The existence of the style-options is not documented in [1].

– If one finds a style-option in a (software) archive, it may need modification to produce the precise effect required.

– It is not obvious how one should refer separately to two items (author and date) supplied via a \bibitem argument originally designed for one. The do-it-yourself-er might have to study style files such as aaai.sty and bbl files such as those produced by aaai-named.bst to deduce how to do this.

• The “short form” scheme seems unsupported.

• At certain points in a document, an author my need to cite several works at once. It may be necessary to specify a page (or section, etc.) for each work. For instance, [3, p. 404] suggests references of the form “(Kelley 1896a, 10; Kelley 1896b; Kelley 1907, 3)”. This is not easy in L^{A}TEX 2.09, since \cite’s optional argument applies to the citation as a whole. The author cannot supply a separate “optional argument” for each work. (By contrast, EndNote’s analogous facilities [33, p. 58] seem to allow each individual work to be given its own “additional text”.)

• It does not seem easy to use different schemes in parallel within the same document. If \cite and thebibliography are defined as required for one scheme, they will usually be unsuitable for any other.

Additional references

As stated in section D.2.2 an author may need to typeset a list of “further reading” etc., in addition to the usual list of “works cited”.

If using the thebibliography environment from one of L^{A}TEX 2.09’s standard styles for such additional references, an author will be faced with the following problems:

• the title will be the same as that for the list of “works cited”, namely “References” for article and “Bibliography” for report and book

• the “labels” (which may be appropriate in the list of “works cited”, particularly for the “reference by number” scheme) will also appear in the additional list (where they are inappropriate), because both lists use the same definition of \bibitem

• by default, the “labels” will not be unique, since the “works cited” list and the “additional references” list will both be numbered from one

• it will be necessary to supply dummy cite-keys, purely to satisfy the syntax required for a \bibitem.

Other problems

The definitions of the thebibliography environment in L^{A}TEX 2.09’s standard styles:

• issue either a \section* or a \chapter* command, using a heading of “References” for article and “Bibliography” for report and book

• set left-head and right-head to either “REFERENCES” or “BIBLI-OGRAPHY”

• do not arrange for a table-of-contents entry.

These definitions can cause problems when the heading, etc. supplied by the standard style is inappropriate, or when a table-of-contents entry is desired. Admittedly, anyone who doesn’t like the standard styles is free to take copies of the style files and modify them to suit their requirements. However, I have the impression that:

• among people who are competent to modify style files, modification of these aspects of the standard styles is “the rule” rather than “the exception”

• those L^{A}TEX users who aren’t particularly computer-literate find the whole business mysterious, and seek out support staff who have to modify these aspects of the style files for them.

Overall, this seems to be an area where L^{A}TEX 2.09 is failing to “free people from formatting concerns to allow them to concentrate on writing” [1, p. 8].

An associated problem is that modified style files may no longer be compatible with standard utilities such as lablst.

D.2.4 Introduction of new facilities

It has been decided [2] that L^{A}TEX 3.0 will be compatible with L^{A}TEX 2.09 input files. Thus, in particular, L^{A}TEX 3.0 must define \cite and thebibliography so that they have the same effect on L^{A}TEX 2.09 input files as the L^{A}TEX 2.09 definitions do. This implies that it would be difficult for L^{A}TEX 3.0 to (for example) define \cite so that there can be an optional argument for each work in a multiple citation and define \bibitem so that it can have an “author” argument and a “date” argument.

It therefore seems best to provide duplicates of the L^{A}TEX 2.09 facilities in L^{A}TEX 3.0 (for “backwards compatibility”) but to attempt to provide new commands/environments in parallel so as to provide the required functionality. The new facilities would be regarded as the “normal” facilities, would be described in the body of the successor to [1], and would be the natural choice for new users. The old facilities would be regarded as “deprecated” and relegated to an appendix of the successor to [1].

Thus we can have both backwards compatibility and improved facilities for the future.

D.2.5 Division of labour

Details needed for document “as a whole”

The three basic citation schemes mentioned in section D.2.2 determine certain details of a document “as a whole”. For each citation, there must be an entry in a reference-list. Each entry in the relevant reference-list must have associated information that can be used in citations.

Details needed for reference-list, etc.

There are a lot of other details that need to be resolved.

The information within each reference-list entry will probably need formatting according to certain rules of “house style”. The information given in citations needs organising in a consistent way (particularly for the “short form” scheme).

Different people may want to assemble their reference-lists in different ways. Some people may wish to “do it themselves” [1, p. 73] from a physical card-index, while some may prefer to use BibTEX to get details from a bib file. In some disciplines, proprietary systems such as EndNote [33] seem popular (because they help the user to search a database for literature to cite, as well as helping the user incorporate details of the literature into a document). Researchers may also wish to incorporate material obtained by searching details held on a cd-rom.

A reference-list generally needs sorting into a particular order. Since the list may occupy several pages, I assume that any sorting is best done outside L^{A}TEX, either by other software (e.g., BibTEX) or manually by the author.

L^{A}TEX and other software

It seems best to regard the movement of text-strings (e.g., an author’s surname) within the “document as a whole” as a task that is distinct from the arrangement of details within the text-strings, and to assume a “division of labour” in which the former task is performed by L^{A}TEX while the latter is performed by some other software or manually by the author. The “division of labour” between L^{A}TEX 2.09 and BibTEX seems to set a good precedent.

This division of labour will lead to modular software. Once the interface between a reference-list and the rest of the document has been defined, people can use L^{A}TEX for the body of their document, but can:

• experiment with different software (BibTEX, EndNote) for formatting the details of their reference-lists

• enhance the other software (e.g., BibTEX) independently of enhancements to L^{A}TEX.

• lay their reference-lists out manually if they prefer.

D.2.6 L^{A}TEX 3.0: A possible user interface?

Specifications and names

If the reasoning given in sections D.2.2, D.2.4 and D.2.5 is accepted, consideration needs to be given to the form that new commands/environments should take in order to support the three basic citation schemes, and to provide facilities for “additional references”. In particular, it will be necessary to choose names other than \cite, thebibliography and \bibitem (since these names will be kept for the facilities provided for compatibility with L^{A}TEX 2.09).

Four sets of commands/environments

It seems to me that it would probably be convenient to have three sets of commands/environments for dealing with citations and the corresponding reference-lists, each set specifically designed to implement a particular citation scheme. Having three such sets gives scope for taking proper account of the peculiarities of each scheme, without having one scheme adversely affected by the peculiarities of another.

To avoid the difficulties mentioned in section D.2.3, it might also be worth having a specific environment for “additional references”.

L^{A}TEX 3.0 might, for example, have commands/environments as specified in the following table.

Citation  Environment for  Entry in
reference-list  reference-list

       Reference by number \numcite        numrefs       \numentry
       Author-date  \dcite         adrefs        \adentry

                             \adcite

       Short form  \firstcite       sfrefs        \sfentry

                             \sfcite

Additional references  —  morerefs      \moreentry

       Analogue at 2.09  \cite     thebibliography    \bibitem

Here it is assumed that:

• \numcite and \numentry have key-list and cite-key (respectively) as their only mandatory arguments.

• \dcite and \adcite have key-list as argument. \dcite gives a citation of the form (1972), while \adcite gives a citation of the form (Crane, 1972). \adentry has three arguments: the cite-key, the author (e.g., Crane), and the date (e.g., 1972).

• \firstcite and \sfcite have key-list as argument. \firstcite gives the form of citation to be used when a work is first mentioned. \sfcite gives the short form to be used in subsequent citations. The arguments of \sfentry might include: the cite-key, the form of reference to be used at the first citation, and the short form to be used subsequently. Whereas \numentry and \adentry can “introduce” the full reference (like \item starts a new item [1, p. 166]), it may be better for \sfentry to have the full reference as an argument, so that it can be used as the default “form to be used at first citation”.

Although it would be desirable for the successors to the “standard styles” to define facilities for all three citation schemes, other \documentstyles need not define facilities for all three. For example, a journal that wants its authors to use the author-date scheme would supply a style file that only provides author-date facilities.

Further details

Reference by number The \numcite and \numentry commands might take the form \numcite{key-list} and \numentry{cite-key}. Notice that, since \numentry is specifically designed for “reference by number”, there is no need to allow an optional label argument like that for \bibitem.

To conform to the ISO specification [28], the successors to the “standard styles” would arrange for \numcite to give a citation of the form (24) and for \numentry to give a reference-list entry of the form

24. Crane, D. Invisible colleges. …

Perhaps L^{A}TEX 3.0 could use the aux file to refine an initial estimate of the width of the final \numentry’s “label”, so that the do-it-yourself-er wouldn’t need to supply a widest-label argument.

Author-date The commands \dcite, \adcite and \adentry might be defined to have the forms \dcite{key-list}, \adcite{key-list} and \adentry{citekey}{author}{date}. The adrefs environment would not have a widest-label argument, since in this scheme entries in the reference-list are unlabelled.

If such a definition of \adentry was documented in the successor to [1], a do-it-yourself-er would be able to use the author-date system just as easily as the “reference by number” system.

To conform to the ISO specification [28], the successors to the “standard styles” would arrange for \dcite to give a citation of the form (1972), for \adcite to give a citation of the form (Crane, 1972), and for \adentry to give a reference-list entry with no label.

Short form The \firstcite, \sfcite and \sfentry commands might be defined as \firstcite{key-list}, \sfcite{key-list} and \sfentry{citekey}[fairly-full-form]{short-form}{full-reference}. Such definitions would, in effect, automate Butcher’s manual method of ensuring consistency [4, p. 178]. Having full-reference as an argument means that the full reference can be used as the default fairly-full-form (to be used when the work is first cited [3, p. 407]). The sfrefs environment would not have a widest-label argument.

The successors to the “standard styles” would arrange for \firstcite to produce the fairly-full-form and \sfcite to produce the short-form.^{1} To conform to the ISO specification [28], neither \firstcite nor \sfcite would add any punctuation (but, for example, a style file that implemented the MLA conventions would have to add brackets [25, ch. 5]). In the successors to the “standard styles”, the \sfentry would produce an entry with no label.

People producing other style files would be free to implement other conventions [4, p. 168].

Some works may have no bibliography or may have just a “select bibliography” [4, p. 168]. For such works, it will still be necessary to supply the details for use by \firstcite and \sfcite. It may therefore be worth allowing a form of sfrefs (e.g., \begin{sfrefs}[null]) that holds details of cited works but does no typesetting.

Additional references Since the entries in an “additional list” will not be cited as such (although an “all sources consulted” list may contain a duplicate of a cited entry in a “works cited” list), the list will be typeset without “labels”. Even in a document that uses the “reference by number” citation scheme (and so needs “labels” in the numrefs list), there will be no “labels” for the entries in an additional list.

1This makes the pessimistic assumption that L^{A}TEX can not itself determine whether a citation is the “first citation” of a particular work. If someone was ingenious enough to produce code that determined whether a citation is a “first citation”, \firstcite would be unnecessary (except, perhaps, for the construction of tables of abbreviations [3, p.

414].)

It therefore seems likely that the morerefs environment could be implemented as a variation of adrefs or sfrefs, the main changes being:

• a change of title (but see section D.2.7)

• absence of cite-key, author, date, fairly-full-form and short-form arguments.

Order within the reference-list

As stated in section D.2.5, it is probably best to leave any sorting of the reference-list to some other software, or to the author.

However, it might be possible for L^{A}TEX to provide a warning if a referencelist is obviously in the wrong order. Perhaps:

• although there may be no easy alternative to numbering \numcites in order of appearance within numrefs (even though “order of first citation” is usually what is required), L^{A}TEX could give a warning if a \numcite gave a number that exceeded the “biggest number produced by \numcite so far” by more than one.

• there could be a warning if an \adentry had an author whose first letter came before the first letter of the preceding \adentry’s author in the alphabet.

Citation of a specific division

As stated in section D.2.3, provision needs to be made for the citation of a particular division (e.g., page, section, chapter, equation) of another work. The syntax of citation commands should not only allow several works to be cited simultaneously, but should also allow the relevant division of each work to be specified.

From the author’s point-of-view, there would be a variety of satisfactory ways to specify citations that are to appear as “[4, p. 10; 5; 6, p. 3]”, e.g.

    \numcite{smith[p. 10],brown,jones[p. 3]}
    \numcite{smith, p. 10; brown; jones, p. 3}
    \numcite{smith & p. 10; brown; jones & p. 3}

The precise syntax would have to take account of the practicalities of programming a command that has to be able to accept pairs of arguments, where the second member of each pair is optional.

Analogous facilities would be needed for author-date and “short form” citations.

Incidentally, since abbreviations such as p., ch., sec., and fig. are common when such divisions are specified, I think that citation commands should arrange for the optional arguments to be typeset with \frenchspacing.

D.2.7 Details of reference-lists

Variations within mainstream publishing practice

Although many academic and technical publications involve only a single undivided reference-list, some such publications involve:

more than one list This situation can arise:

• when there is a list of “further reading” etc., as well as the list of “works cited”. This case has been covered in sections D.2.2, D.2.3, D.2.6 and D.2.6.

• when conference proceedings are produced, since each contribution may have its own reference-list.

• in manuals for software. For example, in [31], a software house’s own publications are introduced in the preface and cited (in effect) using a “short form” scheme, while other people’s publications are listed at the end of the manual and are cited using the author-date scheme.

subdivisions within a list Some reference-lists, particularly in the humanities, are subdivided according to the source of the cited documents (see [4, p. 183], [3, p. 425] and [25, p. 88]).

In some cases, an author may wish to add explanatory paragraphs describing, for example, how material was chosen for a “select bibliography” [3, fig. 15.11] or information about access to (document) archives [3, fig. 15.16].

L^{A}TEX 2.09

L^{A}TEX 2.09 can cope with documents that have more than one thebibliography environment, and seems to deal satisfactorily with a situation in which some \cite commands are to one thebibliography and some are to another (provided that the cite-keys are unique). The default effect is to give labels that are not unique, which will be acceptable when each contribution to a “conference proceedings” has its references numbered from one, but not if “works cited” and “additional references” are both numbered from one (see section D.2.3).

The specification of thebibliography [1, p. 187] does not allow anything other than \bibitems within a thebibliography environment. Hence, it is not clear how one can introduce subheadings within a reference-list. (In practice, a \section* seems to work between \bibitems, but I suspect that it puts L^{A}TEX 2.09 into a loop if placed before the first \bibitem.)

Anyone trying to add explanatory paragraphs (as in [3, fig. 15.11] and [3, fig. 15.16]) will probably find that L^{A}TEX 2.09 objects that

     “Something's wrong--perhaps a missing \item”.

L^{A}TEX 3.0

Ideally, in order to provide support for the conventions that are routine in mainstream publishing practice, L^{A}TEX 3.0 should be able to cope with all the variations outlined in section D.2.7.

Multiple lists The suggestions made in section D.2.6 would probably cater for most situations where a document has more than one referencelist.

The distinction between morerefs and the other environments for referencelists would take care of situations where there is a list of “additional references” as well as a list of “works cited”. The distinction between the “short form” commands/environment and the other commands/environments would take care of situations where a software house uses “short form” for its own publications and some other scheme for other publications. Conference proceedings will be able to have “a reference-list for each contribution” if L^{A}TEX 3.0 follows the L^{A}TEX 2.09 precedent that allows more than one

thebibliography in a document.

Subdivisions and explanatory paragraphs I have the impression that, because the sciences have different conventions from the humanities, people using the “reference by number” citation scheme are unlikely to want the options of subdividing their reference-list and inserting explanatory paragraphs. Therefore, it would be legitimate to say (for example) “subdivisions and explanatory paragraphs are supported within adrefs, sfrefs and morerefs but not within numrefs”, if this made the programming task easier.

For example, it might be convenient to implement numrefs as a “listmaking environment” (as in L^{A}TEX 2.09) but to implement the other environments in some other way. Lack of support for subdivisions and explanatory paragraphs is unlikely to matter in numrefs; the “other way” (more like \paragraph, perhaps?) might make it easier to implement support for these facilities in the other environments.

Other problems One approach to some of the problems mentioned in section D.2.3 is for the standard styles to define the heading for the reference-

list by, for example, \def\numrefsheading{References}, so that anyone

who wants to change the heading can do so by issuing a \renewcommand command somewhere before the start of their reference-list.

Another approach is to work in terms of the standard publishing industry concept of “back matter” [3, p. 4]. Instead of having to understand where the heading, left-head, right-head and table-of-contents entry (or lack of it) originate for units such as the glossary (if any), the reference-list(s) and the index (if any) separately, an author would only have to understand how these features are treated consistently within “back matter”.

Although the “back matter” approach could be used if environments such as numrefs followed the thebibliography precedent and issued commands such as \chapter or \section themselves, authors might find the way that heading, left-head, right-head and table-of-contents entry materialize less mysterious if it was just the same for a reference-list as for (say) a glossary. This would imply that numrefs, adrefs, sfrefs and morerefs should not issue commands like \chapter or \section themselves. As a bonus, it would then be possible for an author to insert an explanatory paragraph before the reference-list, and to arrange for subdivisions.

For example, if there was a backmatter environment within which \chapter was treated as defining a unit of “back matter”, an author’s file might contain commands such as

\begin{backmatter}
   \chapter{Glossary}
      ...

\chapter{References}
\section{Primary sources}
\begin{sfrefs}
...
\end{sfrefs}
\section{Secondary sources}
\begin{sfrefs}
...
\end{sfrefs}
\chapter{Further reading}
\begin{morerefs}
...
\end{morerefs}
\end{backmatter}

D.2.8 Conclusion

L^{A}TEX has a large number of users, and potential users, who wish to produce documents that conform to the conventions that are standard in academic publishing. One element of their requirement is the need to conform to the conventions for citations and reference-lists that are usual in their disciplines.

The choice for L^{A}TEX 3.0 may be between:

1. having more facilities for citations and reference-lists than L^{A}TEX 2.09, perhaps as suggested in section D.2.6. This would imply an increase in:

• the amount of code needed to implement the facilities, and the guru time needed for writing the code

• the number of pages needed, in the successor to [1], to describe the facilities — perhaps 8 pages rather than the 2 pages in [1].

2. no significant increase in the facilities provided for citations and referencelists. Contrary to the idea of “freeing people from formatting concerns to allow them to concentrate on writing” [1, p. 8], many authors (perhaps most authors) would be wasting time:

• hacking at style-files

• searching (software) archives for ready made solutions • taking up support staff’s time in the search for advice (with the support staff in turn taking up gurus’ time in their search for solutions).

Moreover, most of this time would be wasted by (or on behalf of) authors who don’t want anything at all exotic; they just want to conform to the conventions that are standard in traditional academic publishing.

I’m inclined to think that the first option would be the lesser of the two evils.

Appendix: Interface with BibTEX

The preceding sections make some suggestions for a L^{A}TEX 3.0 user interface that would enable the do-it-yourself-er to conform to the conventions that are usual in academic publishing. It is also necessary to consider the implications for the interface to BibTEX.

Single reference-list

Inspecting bst files gives me the impression that, if it was decided to adopt a scheme such as that described in section D.2.6, it would be fairly easy to produce new bst files to supersede existing ones. For example, a bst file that implemented a “reference by number” scheme would write \numentry commands rather than \bibitem commands.

If bst files were created in this way, they would be able to deal with the straightforward situation when there is a single reference-list that contains all works cited (plus, possibly, any works specified by a command like

\nocite).

Multiple reference-lists, all with the same style

More complicated situations can arise in which a document involves several reference-lists. For example:

• The editor of the proceedings of a conference might want the published proceedings to have a reference-list at the end of each chapter.

• If, as suggested in section D.2.7, sfrefs was implemented in a way that allowed a sequence of sfrefs environments to each be preceded by a \section command, then, as far as BibTEX is concerned, each sfrefs environment might be a separate reference-list.

In both these examples, the document would involve several referencelists, but each reference-list would need to be typeset in a common style. I assume that the main problems would be in arranging:

• to have multiple bbl files, or distinct divisions of a single bbl file

• that each reference-list takes its entries from the correct bbl file, or from the correct division of a single bbl file.

Two reference-lists, each with a different style

People producing documents that have a second reference-list (e.g., “further reading”) in addition to the list of “works cited” might want the first list typeset in one style and the second list typeset in another. (In particular, if the “reference by number” scheme is used, the first list will have “labels” but the second list will have no “labels”.) If the commands/environments suggested in section D.2.6 were implemented, the first list would use numrefs, adrefs or sfrefs, while the second list would use morerefs.

In this situation, it would be necessary to communicate to BibTEX that two lists are required, but that they are to be typeset in different styles. Since the second list is to contain “works not cited”, it will also be necessary to specify the works to be shown in the second list.

If it is decided to extend the interface between L^{A}TEX and BibTEX to cater for such situations, it will probably be necessary to consider defining alternatives to L^{A}TEX 2.09’s \bibliography and \bibliographystyle commands, since it seems unlikely that the syntax of the L^{A}TEX 2.09 commands could be extended so as to pass the necessary information. One might, for example, consider syntax such as \bibtexcites[cites-style]{bib-files} and \bibtexmore[more-style]{bib-files}{key-list}. Here, it is assumed that:

• the \documentstyle would set defaults for the cites-style and morestyle that are to be passed to BibTEX, but the user can over-ride the defaults via the optional arguments to \bibtexcites and \bibtexmore

• \bibtexmore’s key-list argument would be used to specify the works to be included in the list of “additional references”.

If it is not practicable to extend the L^{A}TEX/BibTEX interface to cater for these situations automatically, it would presumably be a matter of some “human intervention”:

• to prepare a bbl file for the “additional references”

• to \input the bbl file.

D.3 Further e-mail comments

From:     David Rhead ...
Date:     4 Apr 91 12:16:14
Here are a few comments on Nico's comments (5th March) about my paper on
citations and reference-lists.

>>  2. The paper on reference lists concentrates too much on layout and not
>>  enough on structure. For a, in my humble opinion, much more valuable
>>  discussion of LaTeX 3.0, reference lists and BibTeX I'd like to refer to the
>>  talk Frank Mittelbach gave at the Cork conference last year.
Frank has sent me a copy of "BibTeX reconsidered", by Reinhard Wonneberger
and himself.  I think that this is the written version of the talk that
he gave at Cork.  There seems to be a lot of common ground between my paper
and "BibTeX reconsidered", although the former is looking at it from the LaTeX
point-of-view while the latter is looking at it from the BibTeX point-of-view.
[Both my paper and "BibTeX reconsidered" would like support for multiple
bibliographies, the short title/form scheme (often in footnotes) and
commented bibliographies.]
I'd like "document support" (as Reinhard/Frank put it) for reference-by-number,
author-date and short-form citation schemes.  My paper suggested one possible
user interface through which this support could be provided.  Of course,
other interfaces could be defined that would do the job.  I think that the
main thing is to provide satisfactory support for the 3 schemes.

>>  I'm totally opposed to the idea of having different coding schemes for
>>  different systems of citation. In my opinion, this goes completely against
>>  the basic idea behind LaTeX and SGML, namely separation of form and contents.
>>  Consider the amount of re-coding when switching from the number system to the
>>  name-year system!
There are differences between the schemes that may be so significant that
they could/should be regarded as "different in form".
*   For instance, ISO 690 gives the example
        The notion of an invisible college has been explored in the
        sciences (24).  Its absence among historians is noted by
        Steig (13, p. 556).  It may be as Burchard (8) points out ...
    which, if converted to author-date, would be

        The notion of an invisible college has been explored in the
        sciences (Crane, 1972).  Its absence among historians is noted by
        Steig (1981, p. 556).  It may be as Burchard (1965) points out ...
    where the substitution has to take account of whether the author's name
    does or does not occur naturally in the sentence.  I.e. the form of
    citation depends on what else is in the sentence, and conversion
    cannot easily be automated.  Similarly, I'd be surprised if one can
    guarantee to be able to change from other schemes to the
    short-form scheme without some re-writing.
*   Reference-by-number involves keeping track of one thing (the number),
    author-date involves keeping track of two (surname and date), short-form
    involves keeping track of at least two things (form for first citation,
    form for subsequent citations, perhaps also the full form as the default
    form for first citation).
If the schemes are different in form (i.e. involve different logical
structures), it may be legitimate to consider having different coding
schemes, while still aiming at separation of form from content (e.g.
whether reference-by-number uses superscripts or brackets).
However, if someone has sufficient insight to be able to propose a single user
interface that can cater for all three schemes (plus possible "additional
references") within one set of commands/environments, I agree that it
would be very nice.  My own attempts to define a single interface
that would cater for all three schemes have ended up being unsatisfactory.
Here's how they end up unsatisfactory.
                              ------------
For the purpose of this account I'll use \refentry to mean the successor
to \bibitem.  Presumably \refentry would have to have 2 or 3 arguments besides
the cite-key (rather than \bibitem's one extra argument) so that the arguments
could be used for:
- nothing, in the reference-by-number scheme (since LaTeX would supply
  the numbers)
- author and date, in the author-date scheme
- first-citation-form and subsequent-citation-form, for the short-form
  scheme.
Thus, for reference-by-number, the user (or BibTeX) would supply
\refentry{cite-key}{}{} ...
but for author-date they would supply
\refentry{cite-key}{author}{date} ...
while for short form they would supply
\refentry{cite-key}{fairly-full-form}{short-form} ...
[We're already in trouble.  How do we deal with what appears to be the tendency
in short form for the "full reference" to be the default "fairly full form"?
Do we go for \refentry{cite-key}[fairly-full-form]{short-form}{full reference},
and if so what are the implications for the other 2 schemes?]
Or should one go for
\refentry{cite-key}{author}{date}{fairly-full-form}{short-form} ...

so as to give LaTeX all the information it needs to allow completely
automatic switching between citation schemes, separating form from content
but placing a heavy burden on the user (who would have to provide
at least dummy information for both author-date and short-form
schemes, even if they are only going to use reference-by-number)?
For citation commands one might have \cite and \shortcite (following
precedents in the archives) [but see section 6.5 of my paper, about citation
of a specific devision].  By aiming at author-date, one might be able to
get a .tex file that also worked for reference-by-number.  For example,
if \cite and \shortcite both gave a number when a .sty file implemented
reference-by-number, while \cite gave (author, date) and \shortcite
gave (date) when a .sty file implemented author-date, the following
input might work for both schemes:
    The notion of an invisible college has been explored in the
    sciences \cite{crane-72}.  Its absence among historians is noted by
    Steig \shortcite{steig-81[p. 556]}.  It may be as Burchard
    \shortcite{burchard-65} points out ...
I don't think that the reverse would work, i.e. if you think in terms
of reference-by-number while writing, you'll do things like
    Steig \cite{steig-81[p. 556]}.  It may be as Burchard
which, if converted automatically to author-date, would give
    Steig (Steig, 1981; p. 556).  It may be as Burchard
which gives two occurences of Steig, which isn't right.
For short-form, you might interpret \cite as meaning "the form of citation
used at first citation" and \shortcite as meaning "the form to be used
subsequently".  But if you do this, you'll put your \shortcites in
places that are different from the places you'd put them for author-date
For example, you might have (using an ISO 690 example again)
    ... Steig \cite{steig-81}.  It may be as Burchard
    \cite{burchard-65} points out ... Steig \shortcite{steig-81}
    has further noted ...
for the MLA or
    ... Steig\footnote{\cite{steig-81}} ... Steig\footnote{\shortcite{steig-81}}
for most other publications that use short-form.  [This assumes that the first
of these citations of Steig above is actually the first citation of Steig
in the whole document (or chapter).]  But this would not give you the right
input file for author-date, for which you would want
    ... Steig \shortcite{steig-81}.  It may be as Burchard
Again, you might be able to convert automatically TO reference-by-number,
but it seems unlikely that you could convert automatically FROM
reference-by-number.  The problem might disappear if LaTeX itself could work
out which citations are "first citations".  Then someone can put \cites and
\shortcites in the places required for author-date but choose a .sty file
that gives short-form; the .sty file would ignore the distinction between
\cite and \shortcite and would use the fairly-full-form for "first citation"
and the short-form for subsequent citations.  But can LaTeX work out which

citations are "first citations"?
                              ------------
Although one can try (as above) to imagine a single set of
commands/environments that would support all three schemes, it seems to
lead to problems, and I worry that, in making things right for one scheme,
they'll be made wrong for another.  So I gave up on the idea of a single
set of commands, which was why my paper went for the idea of separate
commands/environments tailored to the 3 specific citation schemes.
But, if someone has more success than I did in imagining how a single set of
commands/environments could be simultaneously compatible with all three schemes,
I'd be interested to see their proposals.
Or should one be less ambitions, perhaps aiming to satisfy people who
want reference-by-number and author-date, and not bothering about people
who want short-form?  Things like \refentry{cite-key}{author}{date},
\cite and \shortcite might work for both reference-by-number and author-date.
(People who want reference-by-number would have to supply author and date
information that they might regard as redundant, unless they go
\refentry{cite-key}{}{} or the syntax is something like
\refentry{cite-key}[author][date].)

>>  To David's review I'd like to add that \bibitem's have no sub-division, at
>>  least not one that is indicated by explicit control sequences (`tags').
>>  Instead, the tagging of \bibitem's is done _outside_LaTeX, which has always
>>  struck me as odd.
In theory, it would be nice if the \bibitem's did have subdivisions.  There
might be subdivisions for the do-it-yourself-er that were analogous to
the fields used by the BibTeX-er.
In practice, to sub-divide the \bibitems would involve deciding what the
subdivisions should be, which leads one into questions that cause difficulty
in BibTeX (e.g., "Should it be address or place-of-publication?",
"Is it really worth having booklet separate?", "Is it really worth
having phdthesis separate from mastersthesis?", "Does volume mean
number-in-series or subdivision-of-book?").  It may not be easy to
answer these questions.  I don't think that it is worth delaying
LaTeX 3.0 while answers are sought.  [One could always return to the
question for LaTeX 4.0, if there was one!]

Appendix E

Some e-mail comments on standard styles

From:     David Rhead ...
Date:     3 Mar 91 17:58:49
At Cork, I think that Frank mentioned the idea of both:
1.  supplying style files that emulate the effect of the present LaTeX 2.09
    "standard styles" (for backwards compatibility)
2.  supplying new "standard style" files.  I think he mentioned having
    analogues of the current article, report and book, plus having
    a "conference proceedings" style.
This note is about (2).
In general, I'd suggest that (2) be done in line with traditional mainstream
publishing practice.  [This seemed to be what Phil Taylor was after when,
at Cork, he described the adjustments he'd had to make to get a LaTeX-ed
book that didn't scream "I've been produced by (La)TeX!", and could be
published.  It would be nice if people in his position didn't have to make so
many adjustments.]  This leaves the problem of determining what mainstream
practice is.  Here are some comments about page-sizes and typefaces.

                           JOURNAL ARTICLES
As regards journal articles, the book by Page, Campbell & Meadows may be
helpful [1, pp. 35-6].  The gist of it seems to be that, for reasons
connected with the sizes of printing presses, journals tend to ignore
B5 and A4 (or did when [1] was written) and go for page sizes of:
*  244 mm x 172 mm, with a 2-column layout.  This is slightly smaller
   than B5 (because B5 doesn't make optimum use of the presses).
   Page et al. seem to be talking in terms of typesize of "9 on 10.5"

   for such a design (although strictly speaking this was for a B5 example.)
   Perhaps this is what the analogue of \documentstyle[9pt,twocolumn]{article}
   should be designed for.
*  276 mm x 219 mm (demi quarto), with a 2-column layout.  This is slighly
   smaller than A4.  Page et al. seem to be talking in terms of typesize of
   "10 on 11.5" for such a design.
   Perhaps this is what the analogue of \documentstyle[10pt,twocolumn]{article}
   should be designed for.
Page et al. also give two B5 (250 mm x 176 mm) examples:
*  single-column, type-area 197 mm x 130 mm, 10 on 11.5
   Perhaps the analogue of \documentstyle[10pt]{article} should
   implement something like this.
*  double-column, type-area 206 mm x 133 mm, 9 on 10.5
These may be more to illustrate economics of different designs than
to say that people actually use B5 much, though.  They say that
single column with 10pt is usually preferred for maths & physics.
If some of the "new standard style" files implemented designs aimed
at the above sizes of paper (with crop marks to show the corners of
the target area), they might serve as a good starting point for anyone who
has to produce a style file for a real journal, particularly if the rest
of the design was based on "the average design" of some real journals
Such styles might also keep authors happy who want to see what their paper
might look like in a real journal.

                                  BOOKS
As regards books, Hugh Williamson suggests that the A series haven't caught
on for book work (not in 1983 Britain, anyway) [2, ch. 3].  He lists the
British Standard cut-page sizes, including (in millimetres):
                              quarto            octavo
crown                        246 x 189        186 x 123
large crown                  258 x 201        198 x 129
demy                         276 x 219        216 x 138
royal                        312 x 237        234 x 156
Apparently 181 x 111 and 178 x 111 are used for paperbacks.
He also lists the corresponding American stock sheet sizes, which give
cut-page sizes of:
      140 x 216
      156 x 235
      127 x 187
      137 x 203
      140 x 210
      143 x 213

which agrees with the Chicago Manual of Style [3, p. 623].
Obviously there is a great variety of sizes.  Presumably any new "standard
book styles" would have to be designed for a particular size of paper.  I'd
suggest that they be designed for a cut-sheet size that IS actually used for
books (e.g. one or more of the above) and that crop-marks be produced to
show the "target area".  Ruari McLean [4, p. 130] says that demi octavo is
one of the most "normal" sizes for books, so perhaps 216 x 138 (or the
nearest US size, 216 x 140) should be one of the "target areas".
As regards section headings -
Williamson [2, p. 163] mentions the scheme:
      level A (i.e., \section) in roman capitals;
      level B in small capitals;
      level C in upper/lower case italic,
      level D in upper/lower case italic, followed by a point, run-in
            with text.
"The Chicago Manual of Style" [3, p. 570] mentions:
      level A caps & small caps (or full caps)
      level B in small capitals
      level C in italics, upper/lower case, followed by a point,
            run in with text.
There seems a fair amount of consensus between these gurus.  Perhaps
elements of these schemes could be combined to give a mainstream design
that isn't going to upset anyone.

                        A4 (and US equivalent)
We have the contradition that:
*  LaTeX is a typeSETTING system.  For most books and journals,
   the typesetting tradition uses paper that is smaller than A4, and
   uses fonts of around 10pt.
*  most LaTeX output comes, at least in the first instance, on A4 (or
   US equivalent) paper from a laserprinter, where A4 etc. is an "office"
   paper size that fits in with the typeWRITING tradition (which usually
   involves fonts of arout 12pt).
So the naive user, seeing something from a "book" \documentstyle on A4
in 10pt, starts to ask "Why doesn't it use all the paper?".  Crop
marks, to indicate the page-size for which the design is intended,
might help:
*  such people to understand why the design doesn't fill A4
*  avoid such people getting an a4.sty out of an archive and going
   \documentstyle[a4]{...} (to get text height/width that purports
   to be "for A4 paper") and then wondering why they've got something
   that is difficult to read
*  help anyone who wants to produce a book to visualise the effect

   that was intended by the designer.
Although it may not be clear which, of the large variety of cut-page
sizes in common use, should provide the target "cut-page sizes" for the
"new standard style's" article and book designs (and their variants),
it does seem fairly clear that:
*  the target "cut-page sizes" should generally be smaller than the A4 sheet
   on which output will initially appear
*  the users should be made aware that the target area is different from A4,
   so they don't ruin the design in their attempts to "make it fit A4".
Obviously, if a design is intended for A4 (as it might be for a
draft article, a report or a thesis), the crop marks would be omitted.
Users would then know that they can use the output just as it comes out
of their laserprinter.

                              GENERAL
I'd be inclined to make Times Roman the normal font for running text
in any "new standard styles" (with a switch somewhere to substitute
Computer Modern for anyone who doesn't have Times Roman, or wants
to get "nearly ready for publication" before they switch from
preview-able Computer Modern to less-easily-previewed Times Roman).
I have the impression that Times Roman is "the default font" for running
text in mainstream publishing.  This would cut out one change that publishers
often seem (rightly or wrongly) to ask people to make who are trying to
typeset a book themselves.  [I have nothing against Computer Modern.
I'm just inclined to "bow to the inevitable".]
I'd be inclined to refrain from providing style-files in situations
where they are unrealistic.  For example, if it is very rare for
real journals and books to use a 12 point typeface for running text,
it may not be worth the effort of supporting \documentstyle[12pt]{article}
and \documentstyle[12pt]{book}.  To support such things may involve
asking the question "What design would one have for a 12pt journal?", to
which the real answer may be "One wouldn't actually have a 12pt journal."
As an alternative to "making the cut-page size depend on the font-size"
(which seems the consequence of the 2.09 way of doing things), it might
be worth considering "making the font-size depend on the target page-size".
E.g. rather than having the user go
      \documentstyle[11pt,twocolumn]{article}
and (having measured distance between the crop-marks) deducing the
size of paper for which the design is intended, it might make more
sense for the user to go
      \documentstyle[...b5]{article}

and get a one-column 10 on 11.5 design intended for B5 paper, or
      \documentstyle[...b5,twocolumn]{article}
and get a two-column 9 on 10.5 design for B5, or
      \documentstyle[...demisemioctavo]{article}
to get an error message along the lines "Sorry.  No single-column
design is available for the demisemioctavo paper-size."  Designs could
then be limited to combinations of paper-size and layout that a real
journal might conceivably use.

                            REFERENCES
[1]   Gillian Page, Robert Campbell & Jack Meadows.
      "Journal publishing: principles and practice", Butterworths, 1987.
      ISBN 0-408-10716-2.
[2]   Hugh Williamson.  "Methods of Book Design", Yale University Press,
      1983. ISBN 0-300-03035-5.
[3]   "The Chicago Manual of Style", Chicago University Press, 1982.
      ISBN 0-226-10390-0.
[4]   Ruari McLean, "The Thames and Hudson Manual of Typography",
      Thames and Hudson, 1980.  ISBN 0-500-68022-1.

—

From:     Sebastian Rahtz <S.P.Q.Rahtz@UK.AC.SOTON.ECS>
Date:     Mon, 4 Mar 91 11:11:40 gmt
... writes:
 > My feeling on the standard document styles (and this is the way
 > that I teach this in my LaTeX classes) is that they DEFINE the
 > structures that appear in a document type, but only give an
 > EXAMPLE of the appearance of those structures as printed. My
yes, fine. a good approach. but it is not very efficient if the
examples are not directly useable. you are talking about people
writing new style files, but 99.9% of the punters have no clue even
where the style files *are* let alone what to do to amend them.
I applaud David Rhead's notes. Lamport states clearly that he
consulted document designers when he created the examples styles;
history seems to show that not many people agree with his style
designers, so lets at least try again and make LaTeX acceptable to a
few more people. Would anyone like to claim that LaTeX's defaults are
acceptable to any publisher they have dealt with? I'd suggest that the
defaults are quite suitable for computer science technical reports;
does not LaTeX aspire to be a professional tool?

 > Rather than trying to do something like say, let's make article
 > look as much like some "standard" appearance for articles (good
 > luck), let's create good structure definitions in our styles,
the two are not opposed
I was interested by David's reflection that
 [12pt]{article}
was contradictory. I suggest that the reason it exists is that people
use `article' for 90% of their daily work (like quick reports on what they
are up to, or class notes) *not* for journal articles. its a misnomer,
IMHO. I have never yet produced a document for use outside this
building that did not require a style different from `article' - are
there *any* journals which would accept it? This is no reflection, of
course, on LL's work! I just think the style designers he talked to
are unrepresentative of the profession.
These discussions often concentrate on headings, by the way. Lets not
forget lists. Maybe David can tell us the ISO standard for vertical
space between items in an enumerated list
PS what puts me off going away and playing with these ideas in sample
styles is a slight fear that the style interface will be very
different from what I have now. After hearing Frank talk about the
concept of an environment stack driven by rules, I have been lying
awake at night trying to decide what I think. If that *is* the model
to adopt, then it affects a lot of the ways one thinks about style
files.

—

From:     David Rhead ...
Date:     4 Mar 91 15:43:11
Sebastian asks whether I can find an ISO standard for vertical space
between items in an enumerated list.
I doubt whether there is such an ISO standard (although we don't have the
standards here for me to browse through).  My impression is that ISO might
pronounce on document-structure and on things like SGML, but that
they are unlikely to pronounce on details of typographic design and
"house style".  I haven't been able to find any particular views expressed
about lists in books on typographic design, either (except for Jan White -
see below).
We do have British Standards for browsing.  Those for theses, manuals
and reports (4821, 4884 and 4811) don't express any views on lists
(although they do seem to generally like arabic numbering).

Jan White devotes pages 88-92 of "Graphic Design for the Electronic Age"
to lists.  As a non-guru, I wouldn't lay lists out like Jan White
(not in a book that has non-indented paragraphs, anyway), since it
leaves the reader unclear about where paragraphs end.  In documents
that have non-indented paragraphs, I wouldn't put extra space between
list items (because I'd want the reader to be able to distinguish
between "list within paragraph" and "list at end of paragraph", and hence
to be able to distinguish one paragraph from the next).  But then Jan White
is a guru and I'm not.
Martin Bryan devotes pages 328-333 of "SGML: an author's guide" to how a
particular "sample DTD" treats lists, but this doesn't answer Sebastian's
query either.  (This particular DTD allows the author to over-ride
"house style" by specifying the numbering sequence to be used.  I was
surprised to see such emphasis given to a feature that over-rides
"house style", since I thought that SGML was intended to help deliver
documents that conform to a "house style".)

—

Date:     Mon, 4 Mar 91 16:29:09 CET
Comments: Originally-From: Don Hosek <DHOSEK@HMCVAX.CLAREMONT.EDU>
My feeling on the standard document styles (and this is the way
that I teach this in my LaTeX classes) is that they DEFINE the
structures that appear in a document type, but only give an
EXAMPLE of the appearance of those structures as printed. My
approach to style design is to, after some preliminaries, input
the base style (report/letter/book/article) which defines the
category of documents that I'm working in.
Rather than trying to do something like say, let's make article
look as much like some "standard" appearance for articles (good
luck), let's create good structure definitions in our styles,
make it easy to adjust the styles with the outline I gave above,
and provide multiple versions of how those classes of documents
could appear.

—

From:     N.POPPELIER@NL.ELSEVIER
Date:     Tue, 5 Mar 91 12:45:21 +0000
I'd like to reply to David Rhead's recent contributions to this list.

1. His summary of page sizes and typefaces for book and journal publishing

is of course interesting in its own right, but its relevance with respect to
the discussion going on this list is not high -- this is not meant as a
personal criticism!
Getting the sizes and typefaces right for a document style for scientific
books or journals always turns out to be the easiest part, in my experience.
LaTeX needs more tools for designing page layouts, font sets -- the new font
selection scheme is already a major improvement -- and section headings, but
summaries of page sizes or section heading schemes are a bit beside the point
here.
A few details:
1.1. As for the use of typeface sizes >10pt: production of camera ready copy
on a larger page frame, using \normalsize = 12pt, followed by photographic
reduction, is normal practice here. `What design would one have for a 12pt
journal?' is not the right question.
1.2. Times Roman can never be the normal font as long as the only font _all_
TeX sites have is Computer Modern. Far more important: the new font selection
scheme combined with the virtual-font mechanism enables you to make a
document-style option for _any_font you like.
1.3. As for Sebastian's comment: there is nothing wrong with using `article',
especially with \baselinestretch > 1.0 and in combination with the 11pt or
12pt option, for producing a preprint of a research paper. In our, i.e.
Elsevier's, case the printed version produced by the author is excellent for
conventional copy-editing. For compuscripts, it doesn't even matter what the
author uses to print his article, since we put in a new document style during
the production stage. Only in a very limited number of cases do we accept the
printed version for actual production of the book or journal.
So `are there *any* journals which would accept it?' is the wrong question,
at least in Elsevier's case.

2. The paper on reference lists concentrates too much on layout and not
enough on structure. For a, in my humble opinion, much more valuable
discussion of LaTeX 3.0, reference lists and BibTeX I'd like to refer to the
talk Frank Mittelbach gave at the Cork conference last year.
As for `standards in academic publishing': there aren't any! At one of the
first meetings of the Dutch TeX Users Group there was a discussion about this
and a staff member of Kluwer Scientific Publishers argued that every
publisher has his own standards. It's the same here at ESP:  every publishing
unit within our company has its own standards, and even though there is
something like an `ESP house style', there is also plenty of variation.

I'm totally opposed to the idea of having different coding schemes for
different systems of citation. In my opinion, this goes completely against
the basic idea behind LaTeX and SGML, namely separation of form and contents.
Consider the amount of re-coding when switching from the number system to the
name-year system!
To David's review I'd like to add that \bibitem's have no sub-division, at
least not one that is indicated by explicit control sequences (`tags').
Instead, the tagging of \bibitem's is done _outside_LaTeX, which has always
struck me as odd.

—

From:     MITTELBACH FRANK <PZF5HZ@EARN.RUIPC1E>
Date:     Thu, 7 Mar 91 12:17:48 CET
Here is an answer from Leslie to Davids mail about sizes
and a few comments of my own.
-----------------------------------------------------------------------------
   he'd had to make to get a LaTeX-ed book that didn't scream "I've been
   produced by (La)TeX!", and could be published.
LaTeX output screams "I've been produced by LaTeX" because it has been
produced with the same standard document style that every other LaTeX
user uses.  As long as there are standard styles, that's going to be
the case.  And that's fine with me.
   journals tend to ignore B5 and A4
When there is a standard journal, there will be a standard journal
style.  Until then, journals who want to typeset using LaTeX will
have to design their own styles.
As long as the printers used by 99.99997% of LaTeX users use either
8-1/2" X 11" or A4 paper, the standard LaTeX styles will be designed to
be printed on that paper.
Similarly, until there is a standard-sized book, the standard LaTeX
"book" style will be for 8-1/2" X 11" or A4 paper, so they can be used
while writing the book.  (This may come as a surprise to some of the
younger members of the TeX community raised during the television age,
but books actually have to be written--a process that takes orders of
magnitude longer than typesetting.) Thus, "book" describes the logical
structures that are handled, not the size and shape of the typeset
output.

Leslie Lamport
-----------------------------------------------------------------------------
I agree with David insofar as the standard styles do not conform
to general practice at least in europe.
But the aim of the standard styles is to provide a layout that
is displaying the contents in a logical way, down to six levels
of sectional units etc.
Or say it in another way, the LaTeX styles will work with any kind
of document that use their tags. This is often not true for
special styles.
We certainly have to try making standard styles that will be usable
by many sites without much adjustments, I agree that at the moment
nearly every user (to my experience in Germany) fiddles around with
the settings because the standard University style is sooooo much
different, but we should keep styles that support documents they
the current styles do.
Adding better support for the two main paper sizes namly Laserprinter
A4 or American size is certainly necessary, and should be part of the
document style since it usually effects quite a few parameter.
The current situation in Europe is not satisfactory with 10 different
A4 styles that all have different problems.

—

From:     "Nelson H.F. Beebe" <beebe@edu.utah.math>
Date:     Mon, 4 Mar 91 12:47:55 MST
...
we should make sure to avoid using style file names longer than 8
characters (PC DOS strikes again).  Even though most TeX
implementations on the PC simply drop the extra characters, and find
the right file, when users employ the truncated name and then port
their files back to other systems, they discover that LaTeX cannot
find their style files, and may have to resort to a TeXpert for help.
...

—

From:     David Rhead ...
Date:     3 Apr 91 11:57:22

You may remember my entry about "standard styles" on March 3rd.
Here are a few comments about the subsequent comments.
I started on March 3rd by recalling that Frank had mentioned (at Cork) the
idea of having both:
1.  style files that emulate the effect of the 2.09 standard styles,
    book, article, report, etc.
2.  some additional style files, analogous to, but different from,
    the present book, article and report, plus a "conference proceedings"
    style.
[I hope I've remembered Frank's remarks correctly, and that (2) isn't just
a figment of my imagination.]
My suggestions were about (2).  I don't know how definite Frank's plans
are about "having some additional style files".  If no-one has time to do
anything, neither my suggestions nor the subsequent comments matter
much anyway.  But if someone has time to do something about (2), they might
as well try to make the style files practically useful.  This could be
either by having designs that are aimed at a phototypesetter, on which
       "print size" will also be "publication size" (which was the scenario
       I assumed)
or     by having designs that are intended for printing "too big" followed
       by photographic reduction for publication (as mentioned by Nico).
If something is done, please could the corresponding style files have
comments stating the design assumptions, e.g. "This design is intended
for ultimate publication (without reduction) on demi-octavo paper.  If printed
on bigger paper, it produces crop marks to show the demi-octavo target area."
or "This design is intended for ultimate publication on B5 paper.  To achieve
the effect intended by the designer, you must photoreduce the LaTeX-ed
output to 70%" or "This design is only intended for use while you are
writing your book.  Unless you are proficient at writing LaTeX style-files,
you are advised to submit your book to a publisher whose staff can supply
style-files that will re-format the book prior to publication.".
I.e. the assumptions should be made clear, so that people know what they
have to do to get the effect the designer intended with the style-files that
form part of the standard distribution, and can make changes (if necessary
to keep their publisher happy) from a position of understanding the
designer's assumptions/intentions rather than from a position of ignorance.
I'd have thought that Leslie's concern about people who are at the "writing"
stage, and are using standard laserprinter paper, would be catered for by
(1), i.e. the style-files that emulate the 2.09 standard styles.
Frank reports Leslie as saying
      "book" describes the logical structures that are handled
This is obviously true in the sense that LaTeX 2.09's book.sty etc. define the
logical structures that LaTeX 2.09's "book style" handles.  But what if

the logical structures that the 2.09 book.sty etc. handle aren't quite
the logical structures of "a book" as understood by the rest of the publishing
industry?  See, for example, pages 4 and 5 of the "Chicago Manual of Style"
(which has its origins pre-television) and pages 157-161 of Jan White's
post-television "Graphic Design for the Electronic Age".  A move towards the
industry's structures would be "a good thing" (e.g. it would make it easier to
implement design decisions like "within front matter we do this" and "within
back matter we do that").  If questions arise about whether any
   "additional style files, analogous to, ... the present book, etc."
should implement 2.09 structures or "publishing-industry standard" structures,
I'd suggest a move towards the "publishing-industry standard" structures.

—

From:     Don Hosek <DHOSEK@EDU.CLAREMONT.HMCVAX>
Date:     Sun, 3 Mar 91 12:23:00 PST
My feeling on the standard document styles (and this is the way
that I teach this in my LaTeX classes) is that they DEFINE the
structures that appear in a document type, but only give an
EXAMPLE of the appearance of those structures as printed. My
approach to style design is to, after some preliminaries, input
the base style (report/letter/book/article) which defines the
category of documents that I'm working in.
Rather than trying to do something like say, let's make article
look as much like some "standard" appearance for articles (good
luck), let's create good structure definitions in our styles,
make it easy to adjust the styles with the outline I gave above,
and provide multiple versions of how those classes of documents
could appear.

Appendix F

Could L^{A}TEX do more for chemists?

This appendix is based on a “paper” that I e-mailed a while back. It also contains some e-mailed comments.

F.1 Motivation

From:     Rainer Schoepf <Schoepf@DE.DBP.ZIB-BERLIN.SC>
Date:     Mon, 15 Oct 90 13:09:23 +0100
I got the following message a while ago, as a reaction to a discussion
I had with John Simmie after the talk at Cork.  He's a chemist, so his
focus of interest is different from the mathematicians one.  I think it's
interesting to pass this on to the list.
 - Rainer
----------------------original message follows ------------------------------
...                    I would like to see ...
a user-definable environment which would parallel displaymath-mode.
For example, in printing chemical equations the conventions are that
the text is in roman not math-italic---ok this is easily done via
\def\chem{\rm\everymath={\rm}}
and \chem can now be included inside $...$ or \[ ... \] or $$ ... $$ etc
or outside as  for {\chem\begin{eqnarray} ... \end{eqnarray}}
Your virtual fonts approach will also work.
Problems arise if mathematical and chemical equations are mixed in the sense

text
text
\begin{equation}      pV = nRT \end{equation}
text
\begin{equation}\chem
                      2 H_2 + 2O_2 = 2H_2O \end{equation}
text
because the automatic numbering schemes assigns (1) to the first eqn
and (2) to the second.  Whereas one wants to number maths eqns and
chemical eqns (reactions) independently---for example,
(2.1) eqn 1 of chapter 2 and (1) reaction number 1, maybe in bold or
oldstyle and maybe with <> brackets.
At the moment the only solution is to hack into latex.tex
and duplicate LL's \equation and \eqnarray environments with new
counters and parameters---this I have done and it works fine for
simple examples but failed for cases where the number of stacked
eqnarray's exceeded the page size (hash size and other mysterious errors).
You suggested that AMSTEX might do but on reflection that involves
even more maths---and what us chemists want is less!
Another feature is the lack of a construct to do the following:
                                catalyst
                    2H_+ O_2 ------------->  2 H_2O
                               300 K; 4 bar
where the length of the arrow (---> or <--- or <===>) is
determined by the amount of text over and under it.
Note that we do not want to disable the equation environment just
have a second customisable one.
Regards, John.

F.2 Typesetting and artwork

It may help to consider the division of labour within a traditional publishing house. A “copy-editor” [4, p. 236] might divide the work on chemical formulae up between “the printer” and “the draughtsman”. Thus, some formulae can be typeset, but others are treated as “artwork” and are drawn.

Although there have been valiant attempts [34, 35] to define TEX macros for drawing chemical structure diagrams, I think it is inevitable that, whatever TEX macros are defined, there will be chemists who come along with requirements that are beyond the abilities of the macro packages. Therefore, it seems sensible to retain the traditional division between typesetting and artwork: to typeset those formulae that can be typeset easily and to get the other formulae drawn in some way.

In TEX terms, artwork can be treated as “graphics” to be “pasted” into a typeset document via \special. For example, chemdraw [36] can produce “encapsulated PostScript”, so a L^{A}TEX-ed document with chemdraw diagrams can be printed on a PostScript printer.

F.3 Desirable facilities for chemists in the new L^{A}TEX

L^{A}TEX is never going to be a system for producing “artwork”, so it seems to me that it would be unwise to attempt to provide comprehensive facilities for “chemical artwork” in L^{A}TEX. However, it does seem worth providing a limited number of facilities to make it easier to produce the chemical formulae that should be treated as “typesetting”.

F.3.1 Environments

Chemists’ problems start when they use L^{A}TEX’s “mathematics” environments for “chemistry”. Chemistry is not mathematics, and the conventions for typesetting chemistry are different from those for typesetting mathematics. In terms of the L^{A}TEX philosophy [1, p. 6] “mathematics” and “chemistry” represent “logically distinct structural elements”.

It would seem natural to:

• define environments for chemists that are analogous to the environments that are available for mathematicians

• within these “chemistry” environments, aim to keep to whatever typesetting conventions are usual in chemistry.

How about defining chem, displaychem and chemequation environments,

by analogy with math, displaymath and equation?

If such environments were defined, the style-file writer would then have control over “mathematics” and “chemistry” separately. In particular:

• The default would be \rm for chemistry (although a designer could change the default in a .sty file). Individual authors would no longer have to search through “double bend” sections of the TEXbook themselves [16, pp. 163 & 179].

• A designer could implement a house-style in which mathematical and chemical equations are numbered in separate sequences [4, p. 224] or a house-style in which there is only one sequence of numbers [37, p. 132].

• It might be possible to arrange that subscripts will normally be at the same level [16, p. 179] inside the environments for chemistry.

F.3.2 Commands

At first sight, the L^{A}TEX manual [1, ch. 3] gives the impression that L^{A}TEX 2.09 provides the “building blocks” to give all the arrows, harpoons and annotation that a chemist could want. However, it is often difficult to get these building blocks arranged in the ways required. For example:

• How does one obtain CH_{3}(C=O)OCl [4, p. 235] in “math mode”? We can’t simply use “=” to mean “double bond”, since TEX puts space around it.

• To represent a reversible reaction with rate constants above/below a pair of harpoons, I ended up with

      \[
      \renewcommand{\arraystretch}{0.5}
      A     \begin{array}{c}
            \scriptstyle k\_1      \\ [1mm]
            \rightleftharpoons     \\
            \scriptstyle k\_{-1}
            \end{array}                 B
      \]

before it looked right. Surely individual L^{A}TEX users shouldn’t have to re-do the “tuning” needed to get these things right?

• As in the above example, arrows and harpoons are often labelled to show reaction conditions. It is not clear how to get arrows/harpoons that expand to the width of the labels.

Many of these difficulties are another consequence of trying to use, for chemistry, the structural elements that were designed for mathematics.

So what commands might usefully be made available inside some future “chemistry” environments?

• It seems desirable [4, p. 237] to have documented facilities for single and double bonds. Triple bonds might also be needed [38]. Documented facilities would also be useful for representing

– single bonds by raised dots [7, p. 59]

– “association of an unspecified type” by three centred dots [37, p. 96].

Might commands such as

\bond, \doublebond, \triplebond and \association

be appropriate?

• It seems desirable to have specific commands for arrows/harpoons with labels above/below (to indicate conditions or rate constants). We can obtain an indication of the combinations of arrows/harpoons that might have traditionally been typeset from [3, p. 371]. Thus, as well as providing simple arrows for one-way reactions, it might be worth aiming to provide commands for: equilibrium reactions (beginning at left and right); reversible reactions (beginning at left and right); reactions beginning at left/right and completed to left/right. Might it be worth defining some commands such as

\oneway, \equilibriumR, \equilibriumL, \reversibleR,
\reversibleL, \rightright, \rightleft, \leftright, \leftleft,

that each accept two parameters: one to give a label above the symbol, the other to give a label below the symbol? For example,

\begin{displaychem}
  2H\_2 + O\_2  \oneway{catalyst}{300 K; 4 bar}  2 H\_2O
\end{displaychem}

might be a natural way to specify But oneway, etc. should ensure that symbol width exceeds label width

2H_{2} + O_{2} ⟶_{300 K; 4 bar}^{catalyst} 2H_{2}O

F.4 Work involved

I’d guess that my suggestions about environments could be implemented by slight modications of the code that implements the corresponding environments for mathematics.

Some new work would be required for commands such as \bond, …,

\association, \oneway, . . . , \leftleft. (The only facility like \oneway,

…, \leftleft that I’ve found in an existing macro package is that for arrows in AMS-TEX [39, p. 140] but, from a chemist’s point-of-view, this doesn’t provide sufficient choice of symbols.)

It might be worth seeking advice about objectives from people who typeset chemistry professionally.

F.5 Documentation

If the facilities that I’ve outlined above were provided, I would be inclined to give them less prominence in the documentation than the analogous facilities for mathematicians. This would help give the impression that, whereas mathematicians can expect L^{A}TEX to do everything they want, chemists can only expect L^{A}TEX to do a certain proportion of what they want. For example, in the successor to [1], the description of facilities for chemists might be relegated to an appendix, which could start with a paragraph explaining that the facilities are intended to support “typesetting” but not “artwork”.

F.6 Some e-mail comments

From:     CA_ROWLEY@UK.AC.OPEN.ACS.VAX
Date:     Mon, 22 Oct 90 13:51:21 GMT
This is a comment on the message from John Simmie, passed on by Rainer:
Whilst I think that the typesetting of chemical formulas, etc, is an
important area for any TeX-based system to provide.  I think that, in
common with maths, it should be provided by an add-on package (as
AMS-LaTeX does for maths).  This should be written in close
cooperation with experts in such typesetting.
Nevertheless, we should attempt to ensure that we dot make any
decisions about "basic LaTeX" which make such an add-on dificult to
produce.  It is not clear to me how we can achieve this in general
(since it applies to many other specialist areas of typeseting)

but I cannot see anything in John's requests that requires any changes
to basic LaTeX.

—

From:     David Rhead ...
Date:     24 Oct 90 14:10:44
%  A LaTeX input file
\documentstyle[11pt]{article}

\begin{document}
%  A chemist might hope to typeset single/double/triple bonds in
%  the following "obvious" ways.  However, none of them will give
%  what is required (because, although the bonds are visually
%  similar to some mathematical sybols, they are logically different).
A single bond? $X-Y$
A single bond? $X.Y$
A single bond? $X\cdot Y$
A double bond? $X=Y$
A triple bond? $X\equiv Y$
\end{document}

—

Subject:  Re: Facilities for chemists.  A LaTeX example file.
From:     bbeeton <BNB@COM.AMS.MATH>
Date:     Wed, 24 Oct 90 18:14:13 CET

in looking at david rhead's chemistry example, i find two distinct
notions presented:
   (a) "the bonds are visually similar to some mathematical symbols"
and
   (b) "they are logically different"
why is it necessary to use the existing notation, just because of (a)?
why isn't it possible, yea almost trivial!, to define new names and
structures for these that are logically distinct?  tex, after all, is

a powerful macro language -- one of its main, perhaps most important,
selling points.
i offer another example to illustrate what i'm suggesting.
   $\vector a$  (or $\vector{a}$ )
assume this will be used only in math mode.
if you're a mathematician this becomes
   \def\vector#1{\vec{#1}}
if you're a physicist, it becomes
   \def\vector#1{{\bf#1}}
here the logical notion is the same, but the representation is different.
seems to me that this is what sgml is trying to tell us.  teach the
user to capture the logic, and let the representation be taken care of
behind the scenes.
i don't say that the available fonts and structures are adequate for
chemistry (they're almost certainly not), but that doesn't mean that
suitable new ones can't be constructed.

—

From:     David Rhead ...
Date:     25 Oct 90 14:08:11
...
... with reference to Barbara's message, there may be no difference
between what she's suggesting and what I'm suggesting.   ...
...                                                    I just intended
to show what a chemist might actually do on seeing the current LaTeX
manual (before the chemist found that it didn't give him/her what s/he
wanted), not to show what I think chemists should do in future.
I agree with Barbara that, in an ideal world (to which we are working with
LaTeX 3.0), the user should not use the existing notation ...

Bibliography

[1] Leslie Lamport. L^{A}TEX: A Document Preparation System. Addison-Wesley, 1986

[2] Frank Mittelbach and Rainer Schöpf. “With L^{A}TEX into the Nineties”, TUGboat, 10 (1989), pp. 681–690.

[3] The Chicago Manual of Style. Thirteenth Edition. University of Chicago Press, 1982.

[4] Judith Butcher. Copy-editing. Second edition. Cambridge University Press, 1981.

[5] Ruari McLean. The Thames and Hudson Manual of Typography. Thames and Hudson, 1980.

[6] Hugh Williamson. Methods of Book Design. Third Edition. Yale University Press, 1983.

[7] Hart’s Rules [for compositors and readers at the University Press, Oxford]. Thirty-ninth edition. Oxford University Press, 1983.

[8] John Miles. Design for Desktop Publishing. Gordon Fraser, 1987.

[9] Jan V. White. Graphic Design for the Electronic Age. Watson-Guptill, 1988.

[10] C. M. Sperberg-McQueen and Lou Burnard (editors). Guidelines for the Encoding and Interchange of Machine-Readable Texts. Draft version 1.1. Chicago, Oxford: Text Encoding Initiative, 1990.

[11] Standard for Electronic Manuscript Preparation and Markup. EPSIG, c/o OCLC, 6565 Frantz Road, Dublin, Ohio 43017–0702, USA.

[12] Author’s Guide to Electronic Manuscript and Markup. EPSIG, c/o OCLC, 6565 Frantz Road, Dublin, Ohio 43017–0702, USA.

[13] Reference Manual on Electronic Manuscript and Markup. EPSIG, c/o OCLC, 6565 Frantz Road, Dublin, Ohio 43017–0702, USA.

[14] Markup of Tabular Material. EPSIG, c/o OCLC, 6565 Frantz Road, Dublin, Ohio 43017–0702, USA.

[15] Markup of Mathemtical Formulae. EPSIG, c/o OCLC, 6565 Frantz Road, Dublin, Ohio 43017–0702, USA.

[16] Donald E. Knuth. The TEXbook. Seventeenth printing. Addison-Wesley, 1990.

[17] Draft for development: Recommendations for the presentation of tables, graphs and charts. DD 52. British Standards Institution.

[18] Frank Mittelbach. “A new implementation of the array and tabular environments”, TUGboat, 9 (1988), pp. 298–313.

[19] Kaare Christian. The UNIX text processing system. Wiley, 1987.

[20] Andrew Trevorrow. DVItoVDU User Guide. C version 1.0 for Unix. University of Adelaide Computing Service, 1986.

[21] Chris Torek. dviselect command specification. University of Maryland.

[22] James Clark. dvitops command specification.

[23] Citing publications by bibliographical references. BS 5605. British Standards Institution, 1978.

[24] Citation of unpublished documents. BS 6371. British Standards Institution, 1983.

[25] Joseph Gibaldi and Walter S. Achert. MLA handbook for writers of research papers. Modern Language Association of America, 1988.

[26] Edward J. Huth. Medical style and format. ISI Press, 1987.

[27] International Committee of Medical Journal Editors. Uniform requirements for manuscripts submitted to biomedical journals. British Medical Journal, 1988, 296, 401–405.

[28] Documentation — bibliographic references — content, form and structure. ISO 690. International Organization for Standardization, 1987.

[29] Maeve O’Connor. Editing scientific books and journals. Pitman Medical, 1978.

[30] Gillian Page, Robert Campbell and Jack Meadows. Journal publishing: principles and practice. Butterworths, 1987.

[31] Marija J. Norušis. SPSS-X Introductory Statistics Guide for Release 3. SPSS Inc., 1988.

[32] Presentation of theses and dissertations. BS 4821. British Standards Institution, 1990.

[33] EndNote: A reference database and bibliography maker. Berkeley: Niles & Associates, Inc., 1989. dviselect command specification. University of Maryland.

[34] Roswitha T. Haas and Kevin C. O’Kane. “Typesetting chemical structure formulae with the text formatter TEX/L^{A}TEX”, Computers and Chemistry, 11 (1987), pp. 251–271.

[35] Michael Ramek. “Chemical structure formulae and x/y diagrams with TEX”. In: Malcolm Clark (editor). TEX: applications, uses, methods, Ellis Horwood, 1990.

[36] chemdraw (a program for the Macintosh). Cambridge, Massachusetts: Cambridge Scientific Computing Inc.

[37] Janet S. Dodd. The ACS Style Guide, American Chemical Society, 1986.

[38] Handbook for Chemical Society Authors, The Chemical Society, 1961.

[39] Michael Spivak. The Joy of TEX, American Mathematical Society, 1986.